### v1

# Beyond Patches: Mining Interpretable Part-Prototypes for Explainable AI

Abstract.
Deep learning has provided considerable advancements for multimedia systems, yet the interpretability of deep models remains a challenge. State-of-the-art post-hoc explainability methods, such as GradCAM, provide visual interpretation based on heatmaps but lack conceptual clarity. Prototype-based approaches, like ProtoPNet and PIPNet, offer a more structured explanation but rely on fixed patches, limiting their robustness and semantic consistency. To address these limitations, a part-prototypical concept mining network (PCMNet) is proposed that dynamically learns interpretable prototypes from meaningful regions. PCMNet clusters prototypes into concept groups, creating semantically grounded explanations without requiring additional annotations. Through a joint process of unsupervised part discovery and concept activation vector extraction, PCMNet effectively captures discriminative concepts and makes interpretable classification decisions. Our extensive experiments comparing PCMNet against state-of-the-art methods on multiple datasets show that it can provide a high level of interpretability, stability, and robustness under clean and occluded scenarios.
xAI, Model Interpretability, Unsupervised Concept Mining
ccs: Computing methodologies  Knowledge representation and reasoning
1. Introduction
Deep learning (DL) has revolutionized computer vision, enabling models to achieve remarkable performance across a range of tasks, from object detection to image classification. Despite these successes, the opaque nature of DL models has raised concerns regarding their interpretability and trustworthiness. In high-stakes domains such as healthcare (Fellous et al., 2019), autonomous driving (Kim and Joe, 2022), and security (Chen et al., 2019; Nauta et al., 2023), understanding why a model makes a certain decision is as critical as the decision itself. Interpretability not only fosters trust but also aids in identifying potential biases, debugging models, and ensuring compliance with ethical and legal standards. Consequently, there is a growing demand for deep vision models that are not only accurate but also capable of explaining their predictions in a clear and comprehensible manner (Saranya and Subhashini, 2023; Vilone and Longo, 2020).
Common post-hoc explainability techniques such as GradCAM (Chattopadhay et al., 2018; Selvaraju et al., 2017a) or ScoreCAM(Wang et al., 2020) try to interpret the model’s decision by reverse-engineering and finding the regions of the input image that contribute most to the output after the training process, as shown in Fig. 2(a). In contrast to such black-box reverse-engineering methods, ante-hoc techniques (Nauta et al., 2023; Chen et al., 2019; Rymarczyk et al., 2021; Nauta et al., 2021; Zhu et al., 2024) make the backbone produce interpretable components to explain its decision. This recognition-by-components theory (Biederman, 1987) describes how humans recognize objects by segmenting them into multiple meaningful concepts. Some ante-hoc methods such as ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023), as shown in Fig. 2(b), segment the spatial deep features extracted by the backbone (before the final pooling layer) into fixed, small patch components and learn a set of prototypes from them. These prototypes, activated by the patches, provide an interpretation of the model’s decision and allow users to identify and compare similar prototypes across different input images. However, these methods have significant limitations. Fixed patch sizes can lead to instability in prototype activation for larger areas, which may fail to focus on primitive semantics. Smaller patches, similarly, may also lose their interpretive capacity due to the limited area they represent. Furthermore, many discriminative and explainable attributes require larger regions to capture their semantic meaning effectively to avoid duplicating the same concept in explanation examples (two similar rows at the bottom of Fig.2(b)). In other words, the regions containing the prototypes can be dynamic and require training.
(a) Heatmap-based
(b) Patched-based
Figure 2. Explaining the decision through examples. (a) Heatmap-based methods see the feature backbones as black-box models, and try to locate regions activated by the most important features. (b) With patch-based methods, the features are decomposed into components extracted by image patches, and an explanation is provided by these components.
Direct application of naïve unsupervised part-discovery or Slot Attention-based methods often falls short in capturing primitive, semantically meaningful concepts from images and fails to provide intuitive explanations of model decisions based on extracted prototypes (van der Klis et al., 2023; Locatello et al., 2020). This limitation arises because part features are typically employed merely to identify similar parts across objects, without enforcing conceptual consistency, while Slot Attention features tend to lack explicit prototypical representations required for interpretable reasoning. The result, therefore, is a reduction in the diversity and a lack of interpretability of the prototypes, hindering the model’s ability to provide meaningful and comprehensive explanations for its decisions.
To address these limitations, this paper introduces an intrinsically interpretable image classifier that mines meaningful and interpretable prototypical concepts from dynamic regions in images to explain the model’s decisions. Our unsupervised concept mining network identifies semantically meaningful components, relying only on class labels without requiring additional part labels or annotations. Our part-prototype concept mining network (PCMNet) detects larger, coherent regions of pixels as meaningful prototype parts. This design allows the components to identify relationships between semantically similar prototypes across instances from different classes. Then, we leverage these prototypes to mine a set of primitive yet discriminative concepts used to explain the model’s reasoning directly. For instance, Figure 2 demonstrates an use case where the prototype heatmaps, derived from our pipeline, shows that the decision-making was based on the ”lofty fenders”, ”shiny metal grille” and ”square shaped lights”, to conclude that the vehicle in the image, is a Dodge Durango.
In the first stage, PCMNet employs an unsupervised part discovery mechanism to learn a set of prototypes using a novel center-clustering loss. In the second stage, the model encodes the deep features into a sparse Concept Activation Vector (CAV) (Kim et al., 2018), which provides a spatially and semantically interpretable explanation of the final decision. Each concept activation is computed as the cosine similarity between part features and class-aware prototype centroids. To ensure that these activated concepts are both primitive and class-discriminative, PCMNet applies a clustering algorithm to the part features corresponding to each class-specific prototype. The center of each resulting cluster then serves as a class-aware prototype centroid. Interpretation of the resulting CAV is straightforward, as it is produced by a sparse, non-negative linear layer that guarantees all concept values remain non-negative.
The contributions of this paper are summarized as follows. (1) PCMNet is presented as a part-prototypical concept mining framework for explaining vision classifiers with primitive, prototypical, and interpretable human-meaningful components. (2) Concepts extracted by PCMNet are leveraged from a broader region than single patches to contain more semantical part-prototype abstraction, similar to the human brain, by using a novel contrastive prototype extraction. (3) Two-level clustering mechanisms are proposed for finding interpretable parts at the pixel level and primitive and intuitive concepts at the feature level. (4) Our experiments show that PCMNet explanation is more robust to occlusions, outperforming state-of-the-art patched-based prototype methods according to classification and explainability metrics.
2. Related Work
(a) Post-hoc or heatmap-based explanation: Heatmap-based explainability methods are a widely used family of post-hoc techniques that visualize which parts of an input image contribute most to a model’s decision. These techniques, often referred to as attribution methods, assign importance scores to different image regions to highlight influential features. Gradient-based methods, such as GradCAM (Selvaraju et al., 2017b) and ScoreCAM (Wang et al., 2020), generate heatmaps by backpropagating gradients concerning input features, revealing class-specific activation regions. While GradCAM produces class-dependent heatmaps, other methods like FullGrad (Srinivas and Fleuret, 2019) are class-agnostic, providing a more general view of model behavior across different outputs. Despite their popularity, gradient-based methods suffer from high sensitivity to noise, which can lead to unreliable and inconsistent heatmaps. To mitigate this, gradient-free CAM techniques, such as ScoreCAM, have been proposed to generate more stable and interpretable explanation maps.
Beyond gradient-based methods, attribution propagation approaches offer an alternative way to decompose model predictions into layer-wise relevance scores. Techniques such as Layer-wise Relevance Propagation (LRP) (Bach et al., 2015) recursively distribute relevance through the network, providing a structured breakdown of model decisions. While initially designed for convolutional neural networks (CNNs), recent adaptations have extended these methods to vision transformers (ViTs) (Chefer et al., 2021), leveraging their self-attention mechanisms for improved interpretability. However, despite their effectiveness in highlighting important image regions, both gradient-based and attribution propagation methods remain fundamentally limited in providing high-level, human-interpretable concepts. Unlike our proposed PCMNet, which explains decisions through part-based prototypes, heatmap-based methods do not inherently capture semantic relationships between features, making them less suitable for interpretable concept-based explanations.
Additionally, since these methods are inherently post-hoc, they do not enable real-time user intervention during test-time inference. This means that once a model makes a prediction, users cannot modify or refine the explanations interactively, limiting their practical use in applications requiring human-in-the-loop decision-making. In contrast, PCMNet introduces an ante-hoc, concept-driven interpretability approach that allows for more direct, structured, and interactive explanations by providing semantically meaningful components rather than relying solely on heatmap visualizations.
(b) Slot attention: Slot Attention has emerged as a powerful mechanism for learning object-centric representations by dynamically assigning feature regions to a fixed number of latent slots (Locatello et al., 2020). Unlike traditional attention mechanisms that distribute importance weights across all features, Slot Attention iteratively refines a small set of slots, each representing a distinct object or meaningful component of an image. This allows models to decompose scenes into structured representations without explicit supervision (Kipf et al., 2022). By leveraging iterative updates with attention-based routing, Slot Attention enables more interpretable feature learning, where each slot encapsulates a semantically meaningful part of the input. This property makes it particularly useful for explainable AI, as it provides a structured decomposition of an image, rather than relying on uninterpretable distributed feature maps (Greff et al., 2020).
Recent works have applied Slot Attention in various domains, such as object discovery, segmentation, and interpretable classification (Singh et al., 2022; Monet et al., 2021). In the context of explainable deep learning, it has been used to learn disentangled representations where each slot corresponds to a distinct semantic concept (Karazija et al., 2021). However, these methods typically focus on object-level decomposition rather than part-based representation. Our PCMNet builds on this idea by using Slot Attention-inspired mechanisms to discover meaningful part prototypes rather than full objects, ensuring better granularity in concept-based explanations. By applying slot-like iterative attention within a contrastive prototype extraction framework, PCMNet refines part-based representations, making them more interpretable and aligned with human perception and reasoning.
(c) Concept-based explanation: Contrary to traditional local feature attribution methods that assess the importance of input-level features, concept-based XAI methods (Dreyer et al., 2024; Chen et al., 2019; Rymarczyk et al., 2021; Bach et al., 2015) analyze the role of latent representations in specific layers of a deep neural network (DNN). These methods explore concepts through various mechanisms, including individual neurons, directional representations such as Concept Activation Vectors (CAVs) (Kim et al., 2018), or broader feature subspaces. Early XAI research primarily focused on understanding how these concepts contribute to global decision-making, identifying the most relevant concepts for a given output class (Dreyer et al., 2024; Bach et al., 2015). Some approaches enhance model interpretability by employing local feature attribution techniques to localize and quantify the importance of concepts for individual predictions, thereby enabling concept-based explanations at the instance level (Nauta et al., 2023, 2021). While these methods provide faithful explanations aligned with the model’s internal reasoning, they are vulnerable to performance degradation under occlusion, as their extracted concepts often rely on limited and highly localized discriminative regions of the input. PCMNet addresses this limitation by part-based prototypical concept reasoning to rely on other discriminative parts that are not occluded.
Additionally, although instance-wise concept-based explanations provide deeper insights, they can also be overwhelming and difficult for stakeholders to process, as hundreds of concepts may need to be analyzed for each instance. To address this, some works highlight the benefits of visualizing local decisions within a global embedding space (Dreyer et al., 2024; Donnelly et al., 2022). With PCMNet, we build upon this idea by introducing part prototypes that encapsulate local concepts. This approach reduces the cognitive load for human interpretation by enabling comparisons between individual (local) prediction and prototypical concepts, making explanations more intuitive and structured.
3. Proposed Prototypical Concept Mining Network
Figure 3. Overall architecture of PCMNet. In the first stage, part prototypes are learned from spatial features extracted by the backbone. In the second stage, prototypes are clustered within each class to identify frequently occurring concepts. The center of these clusters is computed as concept prototypes, and the distance between part prototypes and concept clusters determines the activated concepts for the model’s decision-making. This approach enables interpretable and concept-driven classification.
To explain the decision for a classification problem with a training set $\mathcal{T}$ containing $N$ images $\{(x^{1},y^{1}),...,(x^{N},y^{N}\}\in\mathcal{X}\times\mathcal{Y}$, with $L$ classes, the PCMNet extracts primitive and sparse concepts that are activated and discriminate the object. These concepts are described in a limited set of prototypes that existed for almost all images in the datasets (e.g., lights, fenders, and grilles for cars or heads and wings for birds). In other words, the output of our model is the activated concept vector $z_{i}=[z_{p}^{i}]_{p=1}^{K}$, in which $p$ is the prototype index. Fig. 3 shows the overall overview of our approach.
PCMNet transforms features extracted from the backbone model into explainable and interpretable concepts in three steps. Step 1: The model finds semantically independent, meaningful regions from input images in an unsupervised way by minimizing our novel Marginal Cluster Center loss to learn the prototypes that describe these parts. Then, the part features are disentangled to discriminate the class label. Step 2: Once the part features are learned for the training data, for each part inside all instances of each class, a set of shared elementary characteristics is found by applying a clustering approach to these features. Then, the centroid list, which represents the prototypical concepts, is generated by calculating the center of each cluster. Step 3: Extract the Concept Activation Vector (CAV) for the generated centroid list by computing the similarity of extracted features and each centroid. The CAVs are used to discriminate the class of the input and explain the decision. Also, to avoid losing some discriminative concepts that are not prototypical (i.e., specific and unique for this class and not available for others), we combine the features that are located outside of prototypical maps into the CAVs for classification.
3.1. Step 1: Part-Prototype Discovery
Our model is shown in Fig. 3. The input images are given to a backbone to extract the 3D deep features as $F^{i}\in\mathbb{R}^{D\times w\times h}$ where $D$ is the channel size and $w,h$ denotes the width and height. Based on extracted deep features $F^{i}$, our model finds semantically independent meaningful regions that can be shared in different inputs. We name them “part-prototype”. For determining regions, the model scores each pixel in $F^{i}$ to specify their belongingness to each part-prototype class. The mapping score for each pixel in the location $(q)$ is represented by $M^{i}$, we have $\sum_{p=1}^{K+1}M^{i}_{p}(q)=1$. The spatial feature for each part, $F^{i}_{p}$, is computed by element-wise multiplication of $M^{i}_{p}$ to $F^{i}$. The part features, $\mathbf{f}^{i}_{p}$ is computed as:
(1) $\mathbf{f}^{i}_{p}=\mathcal{GP}(\texttt{conv}(F^{i}_{p}))\in\mathbb{R}^{d_{f}},$
where $\mathcal{GP}$ is the global average pooling and $d_{f}$ is the dimension of features. The last index (K+1) is used to cover the background or non-part regions of the foreground (Alehdaghi et al., 2025). We name these features non-prototypical concepts that are computed as:
(2) $\mathbf{g}^{i}=\mathcal{MP}(\texttt{conv}(F^{i}_{K+1}))\in\mathbb{R}^{d_{f}},$
where $\mathcal{MP}$ is the max-pooling. We use the PDiscoNet (van der Klis et al., 2023) as the baseline for extracting parts features. Our part loss comprises their final loss ($\mathcal{L}_{\text{bl}}$) without their orthogonal component and our marginal cluster center loss ($\mathcal{L}_{\text{mcc}}$):
(3) $\mathcal{L}_{\text{part}}=\mathcal{L}_{\text{bl}}+\alpha\mathcal{L}_{\text{mcc}}.$
The goal of $\mathcal{L}_{\text{mcc}}$ is to ensure the selected region for each part remains consistent across all inputs and serves as a prototype; it must capture similar semantic information across different classes. Additionally, these sub-part features should be both class-descriptive and contain unique ID-aware information.
3.1.1. Marginal Cluster Center Loss
We proposed learnable prototypes as features to be distant from each other and then pushed the model to discover the parts from input images to be close to these prototypes. A soft margin is used to make space for the same part from input images with different classes.
(4) $\mathcal{L}_{\text{mcc}}=\sum_{p=1}^{K}[||\mathbf{f}^{i}_{p}-c_{p}||-m_{1}]_{\star}+\frac{1}{K}\sum_{q=p+1}^{K}[m_{2}-||c_{p}-c_{q}||]_{\star},$
where $c_{p}$ is a learnable prototype for part $p$ and $[\cdot]_{\star}$ is $\max(\cdot,0)$.
3.2. Step 2: Part-Centroid List Generation
During training in Step 1, the backbone learns to extract and use part-prototype features from the input data to fulfill its training task. To find all relevant spatial and class-aware concepts, as shown at the top of Fig. 4, we collect all extracted features from the backbone for each class and each part and then cluster them to have a minimal set of plain concepts that consistently describe the visual attributes of those parts. Multiple concepts are considered. They can be seen in each part of each class for more complex objects. To mine those concept prototypes, the centroids of each cluster are computed as $C(l,p,j)\in\mathbb{R}^{d}_{f}$ for a clustering index of $l$, part $p$, and a class of $j$.
3.3. Step 3: Concept Activation Mining
Once the centroid list is generated, each concept’s activation value is defined by the similarity of extracted features to each centroid:
(5) $\mathbf{z}^{i}=[\mathcal{S}(C(l,p,j),\mathbf{f}_{p}^{i})\;\forall p\in\{1..K\}\;\text{and}\;j\in\{1..L\}]\in\mathbb{R}^{d_{c}}$
where $\mathcal{S}$ is cosine distance and [.][.][ . ] represents the vector concatenation. $d_{c}$ is the number of total concept centroids. We aim to explain each output class of our model using a small set of interpretable concepts. We, therefore, train a sparse classifier on top of $z^{i}$ to obtain the final classification scores $o^{i}=W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b$ and the predicted class $\hat{y}^{i}=\arg\max(o^{i})$. Here, $W_{1}\in\mathbb{R}^{d_{c}\times L}$, $W_{2}\in\mathbb{R}^{d_{f}\times L}$ and $b\in\mathbb{R}^{L}$ denote the classification weights and bias term, respectively. This layer is trained with the following classification loss, using the GLM-SAGA optimizer (Wong et al., 2021):
(6) $\mathcal{L}_{\text{cl}}=\mathcal{L}_{ce}(W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b,y^{i})+\lambda\mathcal{R}(W_{1}),$
where $\mathcal{L}_{\text{ce}}$ is the cross-entropy loss, $y^{i}$ is the label, $\lambda$ is the sparsity regularization strength and $\mathcal{R}(W)=(1-\gamma)\frac{1}{2}\|W\|_{F}+\gamma\|W\|_{1}$, here $\|W\|_{F}$ is the Forbenius norm and $\|W\|_{1}$ is the element-wise matrix.
Figure 4. Concept Mining Module. First, a part centroid list is generated by applying DBSCAN(Ester et al., 1996) clustering within each class and computing the centers of the resulting clusters. These centroids serve as frequently occurring concept prototypes. Next, to activate the most relevant concept for a given part feature, the similarity (inverse distance) to the centroids is measured. These values are then used as CAVs to inform the model’s decision-making process .
3.4. Overall Training
We jointly optimize the network in an end-to-end manner by loss:
(7) $\mathcal{L}=\mathcal{L}_{\text{part}}+\beta\mathcal{L}_{\text{cl}},$
where $\beta$ is hyperparameter. For the first step of training, we set $\beta=0$; then, once the part-prototypical concepts are learned, we set $\beta=2$ based on our hyperparameter analysis.
4. Experiments and Results
Method Consistency (Intra) $\uparrow$ Consistency (Inter) $\downarrow$ F(1)-F(5) $\uparrow$ Sp $\uparrow$ Stability C Acc $\uparrow$
Stanford Cars Resnet50+Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 0.16 - 0.23 22.74 65.4 84.1
Resnet50+W $\times$ Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 1.54 - 6.61 22.74 65.4 84.1
ProtoPNet + Act 45.70$\pm$4.9 11.22$\pm$5.5 2.2-55.64 60.92 69.6 84.5
ProtoPNet + W $\times$ Act 45.70 $\pm$ 4.9 11.22$\pm$5.5 9.3-80.12 60.92 69.6 84.5
PiPNet(Covx) + Act 52.78$\pm$7.4 9.07$\pm$9.4 0.02-60.62 70.59 65.2 88.5
PiPNet(Covx) + W $\times$ Act 52.78 $\pm$ 7.4 9.07$\pm$9.4 10.83-95.66 70.59 65.2 88.5
PiPNet(Res50) + Act 42.40$\pm$17.5 17.4$\pm$2.4 0.18 - 1.21 63.19 60.8 86.46
PiPNet(Res50) + W $\times$ Act 42.40$\pm$17.5 17.4$\pm$2.4 9.43 - 93.15 63.19 60.8 86.46
Ours + Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 37.94 - 63.3 57.45 71.5 88.7
Ours + W $\times$ Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 59.33 - 91.73 57.45 71.5 88.7
Birds (CUB 200) Resnet50+Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 0.35 - 1.12 24.14 60.3 81.2
Resnet50+Rel $\times$ Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 2.43 - 5.73 24.14 60.3 81.2
ProtoPNet + Act 51.47$\pm$5.8 10.16$\pm$9.2 1.15-51.96 79.51 63.6 81.45
ProtoPNet + W $\times$ Act 51.47 $\pm$ 5.8 10.16$\pm$9.2 9.36-75.50 79.51 63.6 81.45
PiPNet(Covx) + Act 54.80$\pm$8.3 7.77$\pm$8.0 0.01-54.14 80.72 64.7 85.0
PiPNet(Covx) + W $\times$ Act 54.80 $\pm$ 8.3 7.77$\pm$8.0 5.13-92.83 80.72 64.7 85.0
PiPNet(Res50) + Act 42.28$\pm$3.2 15.57$\pm$15.7 0.14 - 2.35 77.83 65.9 82.0
PiPNet(Res50) + W $\times$ Act 42.28$\pm$3.2 15.57$\pm$15.7 9.40 - 89.77 77.83 65.9 82.0
PCMNet (Ours) + Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 41.82 - 75.24 57.45 67.1 84.7
PCMNet (Ours) + W $\times$ Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 55.13 - 89.88 57.45 67.8 84.7
Table 1. Performance of PCMNet according to xAI metrics on the Stanford Cars and Birds datasets.
PCMNet is validated on standard benchmarks in prototype literature: Stanford Cars (Krause et al., 2013) (196 classes of different make and models) and CUB-200-2011 (Wah et al., 2011) (200 bird species, including attribute annotation for visual features). The ResNet50 model pretrained on ImageNet (Deng et al., 2009) is used as the feature backbone. Each training batch (32 images) is resized to 488×488, randomly cropped, erased, and padded. Optimization follows ADAM (Kingma and Ba, 2014) with a linear warm-up. We set regularization parameters $\lambda$ and $\gamma$ as same in (Oikarinen et al., 2023). $\alpha=1.5$, $\beta=2$, with $m_{1}=0.3$ and $m_{2}=1.5$ chosen via hyperparameter analysis in the suppl. material. More details on the experimental protocol are presented in the supplementary materials.
4.1. Explainability Metrics
To evaluate the explainability of final features, we rely on performance metrics established in the literature (Dreyer et al., 2024) with some minor modifications to express better comprehension.
(a) Faithfulness:: Faithfulness(Dasgupta et al., 2022; Chan et al., 2022) evaluates how much the activated concepts influence the model’s final prediction, serving as a key metric for assessing the quality of explanations. One widely used approach for measuring faithfulness is concept deletion, which involves removing selected concepts from the model’s reasoning process and observing the resulting change in output confidence. Prior works, such as (Dreyer et al., 2024), typically assess this by removing only the single most important concept. However, this limited scope fails to capture the broader contribution of other activated concepts and may not reveal the full extent of their impact. To provide a more comprehensive assessment, we extend this approach by successively removing the top-$k$ most activated concepts and measuring the resulting drop in model confidence. A larger drop indicates higher faithfulness, as it suggests that the removed concepts were indeed critical to the model’s decision. This extended evaluation allows us to better quantify how much the explanation aligns with the model’s internal reasoning.
(b) Stability: The extracted concepts from input images that share some semantic concepts must be stable and explainable for unseen images. To evaluate stability, we compute CAVs on $k$-fold subsets of the data ($k$ = 10 as default) similar to (Dreyer et al., 2024) for all classes at steps 2 and 3. We then map CAVs together using a Hungarian loss function and measure the cosine similarity between them.
(c) Consistency: Activated concepts from images with the same class should be similar to be consistent and explain the model’s decision. To measure this consistency, we extract the CAV from images and then measure the cosine similarity between them. The mean cosine similarity is assessed between images from the same class and between images from different classes. The ratio between inter/intra-class similarity should be high for a stable explanation.
(d) Sparseness: The sparseness metric essentially describes the uniformity of the concept activation, where having some concepts activate more than others is considered easier to interpret. This is because a uniform distribution of concept activations would provide little information on the importance of specific concepts or parts of the images, i.e., a high entropy in the generated concepts.
4.2. Comparison to Prototype xAI Methods
To evaluate PCMNet, we report the classification accuracy and the XAI evaluation criteria discussed in Sec 4 and compare them to other ante-hoc methods ( ProtoPNet(Chen et al., 2019) and PIPNet(Nauta et al., 2023)) in tables 1. The consistency metric is reported in two ways: Intra and Inter, which are cosine similarities between activated concepts of the same class and different classes, respectively. Faithfulness is reported as F(n), measuring accuracy drops when the top n important concepts are removed from the decision. To compute the contribution of each concept or feature to the final decision, we evaluate two strategies: (1) using raw values of each concept (”Act”), and (2) computing a weighted contribution ( gradient) as the product of raw values and its corresponding weight in the classification layer (“W $\times$ Act”).
4.3. Ablation Studies
4.3.1. Number of Prototypes
We assess the impact of the number of parts on both classification accuracy and faithfulness in Table 2. As the model extracts more prototypical concepts, its decisions become more aligned with the underlying features, enhancing interpretability. Interestingly, when the number of prototypes is smaller, the faithfulness scores for the top concepts tend to be higher. This is because the model relies on a more compact set of prototypes, concentrating its decisions on fewer, more influential concepts.
K value Cars Birds
C Acc F(1)-F(5) C Acc F(1)-F(5)
0 (Resnet50) 84.1 1.54 - 6.61 83.2 1.21-5.56
4 83.5 79.96 - 91.01 83.5 71.42-85.64
5 84.1 70.11 – 91.33 84.9 66.7 - 85.9
6 85.2 66.02 – 90.65 85.7 59.9 - 86.7
7 86.7 62.47 – 91.25 86.0 55.7 - 87.3
8 87.5 59.33 – 91.73 85.1 53.0 - 86.4
9 86.4 55.10 – 91.52 84.4 52.5-86.1
Table 2. Impact on performance of the number of prototypes.
Figure 5. Visualizing activated concepts in the test set and examples from the training set for Stanford Cars and CUB datasets.
4.3.2. Effect of Modules
To evaluate the effectiveness of each component in PCMNet, we conduct an ablation study on the baseline PDiscoNet model, which extracts local parts without explicit concept learning. As shown in Table 3, we analyze classification accuracy (C Acc) and feature faithfulness under occlusion (F(3)) on both the Cars and Birds datasets, alongside the model size and FLOPs. Introducing the Marginal Clustering Center (MCC) loss into the baseline improves classification performance (from 84.2% to 86.1% on Cars) and more than doubles faithfulness scores, indicating that enforcing clustering consistency across part features supports more explainability. The inclusion of the Concept Mining Module (CMM)—responsible for extracting sparse and interpretable concepts—dramatically boosts faithfulness (from 4.9% to 55.6% on Birds), even with minimal impact on accuracy. When both MCC and CMM are combined, PCMNet achieves the best results across all metrics, including 87.5% classification accuracy and 67.4% faithfulness on Cars, with only a marginal increase in parameter count and computational cost. These findings underscore the complementary roles of MCC in enforcing cluster-level consistency and CMM in mining high-quality, human-aligned concepts.
Settings Cars Birds #Params FLOPs
C Acc F(3) C Acc F(3)
Baseline(van der Klis et al., 2023) 81.2 5.4 82.3 4.9 27.1M 17.2G
Base + MCC 86.1 9.6 84.5 9.1 27.2M 17.2G
Base + CMM 84.3 59.2 82.1 55.6 27.4M 17.5G
Base + MCC + CMM 87.5 67.4 84.7 64.9 27.5M 17.5G
PipPNet(Resnet) 86.4 62.7 82.0 58.6 25.1M 27.6G
PipPNet(Covx) 88.5 60.6 85.0 57.5 29.3M 32.3G
ProtoPNet 84.5 38.6 81.4 31.7 28.9M 18.4G
Table 3. Classification accuracy and faithfulness using different PCMNet modules.
Figure 6. t-SNE visualization of concept-level part features extracted by PCMNet. Colors represent different prototypes, while shapes indicate object classes.
4.4. Robustness to Occlusion
To evaluate the robustness of PCMNet’s decision-making under occlusion, we conducted controlled experiments in which regions corresponding to the most activated concepts were systematically masked, and the resulting images were re-evaluated by the model. This experiment aims to assess the stability of model predictions when key interpretable regions are removed. Specifically, we identify the center of the most activated concept—based on the model’s response to clean images—using the concept representation (patch for PIPNet and ProtoPNet, mask for PCMNet), and then occlude a rectangular region centered around this point. We consider three levels of occlusion, masking 10%, 20%, and 30% of the image area. PCMNet’s performance under these conditions is compared against two other interpretable baselines: ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023).
Figure 7. Classification Accuracy and Faithfulness under Different Occlusion Levels.
Fig. 7 shows that PCMNet maintains stable accuracy at lower occlusion levels and degrades more gradually than ProtoPNet and PIPNet. This resilience stems from PCMNet’s broader part-based concepts, which introduce redundancy and enable accurate predictions even when key components are missing. Unlike patch-based models that rely on fixed, small regions, it extracts prototypes from semantically meaningful, wider regions. As a result, such methods experience sharp performance drops under occlusion due to their dependence on localized patterns, whereas PCMNet remains more robust by leveraging multiple related part concepts, preventing performance from relying too heavily on a single visual cue.
4.5. Qualitative Results
4.5.1. Concept Activation and Occlusion Robustness
To further illustrate the interpretability and robustness of PCMNet, we compare the activated concepts of our method with PIPNet (Nauta et al., 2023) under clean and occluded conditions, as visualized in Fig. 5. For each method, we select two activated concepts on test images and retrieve training set examples with the highest activation values for those concepts. This comparison shows several key advantages of PCMNet:
Broader Concept Coverage: PCMNet extracts more diverse and semantically rich concepts compared to PIPNet. Our method leverages attention-based masks instead of fixed patches to capture part-level semantics over a broader spatial context, leading to more varied and interpretable regions across different object parts.
More Informative Examples: The training examples retrieved by PCMNet for each concept exhibit more clearly activated and semantically aligned regions. This suggests that the discovered concepts better reflect the underlying factors influencing the model’s decision, thereby enhancing explanation quality.
Robustness to Occlusion: Under occlusion, PIPNet often fails to highlight meaningful or consistent regions due to its reliance on limited regions. In contrast, PCMNet maintains interpretability by activating alternative, unoccluded parts that belong to the same or related concepts. This demonstrates the model’s ability to generalize concept reasoning even when certain visual cues are missing.
4.5.2. Visualization of Concept Prototypes
To better understand the behavior of the learned prototypes in PCMNet, we visualize the extracted part-level features using t-SNE (van der Maaten and Hinton, 2008), as shown in Figure 6. Each point corresponds to a feature from a localized part, colors indicating the prototype (i.e., concept) and marker shapes denoting the object class. This visualization represents several insights:
Disentanglement of Prototypes: Features assigned to different prototypes form well-separated clusters, suggesting that PCMNet effectively disentangles the latent interpretable part-based concept space. This separation implies that the Marginal Clustering Center loss and Concept Mining Module (CMM) encourages diverse and semantically distinct part concepts, which supports both interpretability of model decisions and robustness on occlusion.
Multi-Class Concept Coverage: Many prototype clusters contain the same semantic parts from multiple object classes. This indicates PCMNet learns class-agnostic concepts such as lights, wheels, or windows, which are shared across different categories. Such generalization is crucial for explainability, as it facilitates human-aligned, semantically meaningful concepts that transcend dataset labels.
Intra-Class Concept Diversity: Interestingly, features from the same class are distributed across multiple prototype clusters. This highlights PCMNet’s ability to capture intra-class variation through multiple part-based concepts. For instance, different views or structural variations of a car class may be assigned to distinct concepts (Backlight vs headlight), further enhancing the model’s fine-grained interpretability based on human perception and explanation.
5. Conclusion
We presented PCMNet, a novel framework designed to bridge the gap between accuracy and interpretability in deep learning models. By mining part-prototypes from dynamic image regions, PCMNet improves upon existing patch-based prototype models by ensuring semantic coherence across instances. Our results show that PCMNet enhances faithfulness, concept stability, and explainability, outperforming existing interpretable models such as ProtoPNet and PIPNet. Moreover, our occlusion-based robustness analysis confirmed that PCMNet remains stable under missing or distorted input regions, demonstrating its reliability in real-world scenarios. Unlike conventional post-hoc explainability methods, PCMNet allows for more human-aligned, concept-based reasoning, making deep learning decisions more transparent and interpretable.
Appendix A Additional Experiment and Results
A.1. Implementation Details
We used Resnet50 as the backbone, which extracts 2048 features. We project these features into 512 features before global pooling, i.e., $d_{f}=512$. Based on our experiment and set of hyper-parameters, for the CUBs dataset, the $d_{c}=2954$ and Stanford Cars $d_{c}=3092$ are chosen. As the DBSCAN is used for clustering inside of each class, for each part and class, the number of clusters is not fixed.
The $\mathcal{L}_{\text{bl}}$(van der Klis et al., 2023) loss also is
(8) $\mathcal{L}_{\text{bl}}=\mathcal{L}_{\text{cond}}+\mathcal{L}_{\text{equi}}+\mathcal{L}_{\text{pres}},$
where the objectives for each loss are:
  • $\mathcal{L}_{cond}$: Each detected part should consist of a condensed and contiguous image region.
  • $\mathcal{L}_{equi}$: encourages the equivariance of the attention maps in for input images and their augmented version (translation, rotation or scaling).
  • $\mathcal{L}_{pres}$: All parts should be present at least in some of the images in the batch to avoid finding non-relevant parts.
A.2. Additional Ablation Studies
A.2.1. Effect of Hyper-parameters
We investigate the sensitivity of our model to the hyper-parameters $\alpha$ and $\beta$, which control the strength of the concept consistency and interpretability regularization terms, respectively. As shown in Fig. 8, our model remains robust across a wide range of values. Specifically, the classification accuracy (C Acc) and faithfulness score F(n) exhibit stable trends, with peak performance achieved around $\alpha=1.5$ and $\beta=2.0$. This indicates that PCMNet effectively balances accuracy and interpretability without requiring extensive hyper-parameter tuning. The consistency of the scores across different settings highlights the reliability of the proposed regularization components.
Figure 8. Impact of hyper-parameters $\alpha$ and $\beta$ on classification accuracy (C Acc) and faithfulness score F(3). Our method maintains strong performance across a wide range of values, showing robustness and stability in both predictive accuracy and interpretability.
Also, in Fig. 9, we evaluate the impact of hyperparameters in Eq.4 on classification accuracy. At first, we set $m_{2}=1$ and measure the effect of $m_{1}$ by gradually increasing the values. Once it is found, we search for $m_{2}$. The best performance is achieved when $m_{1}$ is set to 0.3 and $m_{2}$ to 1.5, respectively. The upward trend of the bars demonstrates the effectiveness of each loss.
(a) $m1$.
(b) $m2$
Figure 9. The model’s classification accuracy in different values of $m1$ and $m2$.
Figure 10. Visualization of activated concepts in test data and showing some examples in the training set for StanfordCars datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 11. Visualization of activated concepts in test data and showing some examples in the training set for CUB-200 Birds datasets. Each concept is shown with an index and percentage of explanation for the model decision.
Figure 12. Visualization of non-prototypical activated concepts in test data and showing some examples in the training set for (left) Stanford Cars and (right) CUB-200 Birds datasets. Similar to ProtoPNet and PIPNet, we select the patch for the region of the activated concept.
A.2.2. Effect of Non-Prototypical Concepts
To evaluate the impact of different concept vector types on model performance and interpretability, we conduct an ablation study using prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors. As shown in Table 4, using only non-prototypical vectors significantly reduces both classification accuracy and faithfulness scores. In contrast, employing only prototypical vectors retains higher classification performance and strong concept-based faithfulness. The full PCMNet, which combines both concept types, achieves the best results across both datasets, suggesting that the interplay between prototypical and non-prototypical representations enhances both prediction accuracy and interpretability.
CAVs Cars Birds
C Acc F(1)-F(5) C Acc F(1)-F(5)
Baseline 84.1 1.54 - 6.61 83.2 1.21-5.56
PCMNet + Prototypical 86.2 61.8 – 92.63 84.5 56.8 - 87.4
PCMNet + Non-Prototypical 79.4 42.80 – 75.73 78.15 35.9 - 67.7
PCMNet (Full) 87.5 59.3391.73 86.0 55.7 - 87.3
Table 4. Impact of prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors on PCMNet’s classification accuracy (C Acc) and faithfulness scores (F(1)-F(5)) on Cars and Birds datasets. The full model achieves superior performance, highlighting the benefit of combining both concept types.
A.3. Additional Qualitative Results
More results of the PCMNet prototypical concept are shown in Figs. 10 and 11. 12 illustrates examples of non-prototypical concepts discovered by our model for the Cars (left) and CUBs (right) datasets. For each dataset, the concept location on a test image is highlighted with a green bounding box, while corresponding activations in training examples are shown on the right using red bounding boxes. These concepts provide complementary cues to prototypical ones. For example, in the first row of Cars, the ”cargo” attributes are not well represented in prototypical concepts since they appear in a small subset of vehicles, yet they are highly discriminative for decision-making. Similarly, the last row in CUBs shows the ”webbed feet” concept, which is especially relevant and explainable for Albatross species, while foot prototypes are not sufficiently distinctive across the dataset. These findings highlight the value of incorporating non-prototypical concepts in improving both interpretability and robustness of the model.

### v2

@@
 Beyond Patches: Mining Interpretable Part-Prototypes for Explainable AI
-Abstract.
-Deep learning has provided considerable advancements for multimedia systems, yet the interpretability of deep models remains a challenge.
-State-of-the-art post-hoc explainability methods, such as GradCAM, provide visual interpretation based on heatmaps but lack conceptual clarity.
-Prototype-based approaches, like ProtoPNet and PIPNet, offer a more structured explanation but rely on fixed patches, limiting their robustness and semantic consistency.
-To address these limitations, a part-prototypical concept mining network (PCMNet) is proposed that dynamically learns interpretable prototypes from meaningful regions.
-PCMNet clusters prototypes into concept groups, creating semantically grounded explanations without requiring additional annotations.
-Through a joint process of unsupervised part discovery and concept activation vector extraction, PCMNet effectively captures discriminative concepts and makes interpretable classification decisions.
-Our extensive experiments comparing PCMNet against state-of-the-art methods on multiple datasets show that it can provide a high level of interpretability, stability, and robustness under clean and occluded scenarios.
-xAI, Model Interpretability, Unsupervised Concept Mining
-ccs: Computing methodologies  Knowledge representation and reasoning
-1. Introduction
-Deep learning (DL) has revolutionized computer vision, enabling models to achieve remarkable performance across a range of tasks, from object detection to image classification.
-Despite these successes, the opaque nature of DL models has raised concerns regarding their interpretability and trustworthiness.
-In high-stakes domains such as healthcare (Fellous et al., 2019), autonomous driving (Kim and Joe, 2022), and security (Chen et al., 2019; Nauta et al., 2023), understanding why a model makes a certain decision is as critical as the decision itself.
-Interpretability not only fosters trust but also aids in identifying potential biases, debugging models, and ensuring compliance with ethical and legal standards.
-Consequently, there is a growing demand for deep vision models that are not only accurate but also capable of explaining their predictions in a clear and comprehensible manner (Saranya and Subhashini, 2023; Vilone and Longo, 2020).
-Common post-hoc explainability techniques such as GradCAM (Chattopadhay et al., 2018; Selvaraju et al., 2017a) or ScoreCAM(Wang et al., 2020) try to interpret the model’s decision by reverse-engineering and finding the regions of the input image that contribute most to the output after the training process, as shown in Fig.
-2(a).
-In contrast to such black-box reverse-engineering methods, ante-hoc techniques (Nauta et al., 2023; Chen et al., 2019; Rymarczyk et al., 2021; Nauta et al., 2021; Zhu et al., 2024) make the backbone produce interpretable components to explain its decision.
-This recognition-by-components theory (Biederman, 1987) describes how humans recognize objects by segmenting them into multiple meaningful concepts.
-Some ante-hoc methods such as ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023), as shown in Fig.
-2(b), segment the spatial deep features extracted by the backbone (before the final pooling layer) into fixed, small patch components and learn a set of prototypes from them.
-These prototypes, activated by the patches, provide an interpretation of the model’s decision and allow users to identify and compare similar prototypes across different input images.
-However, these methods have significant limitations.
-Fixed patch sizes can lead to instability in prototype activation for larger areas, which may fail to focus on primitive semantics.
-Smaller patches, similarly, may also lose their interpretive capacity due to the limited area they represent.
-Furthermore, many discriminative and explainable attributes require larger regions to capture their semantic meaning effectively to avoid duplicating the same concept in explanation examples (two similar rows at the bottom of Fig.2(b)).
-In other words, the regions containing the prototypes can be dynamic and require training.
-(a) Heatmap-based
-(b) Patched-based
-Figure 2.
-Explaining the decision through examples.
+Abstract
+As AI systems grow more capable, it becomes increasingly important that their decisions remain understandable and aligned with human expectations.
+A key challenge is the limited interpretability of deep models.
+Post-hoc methods like GradCAM offer heatmaps but provide limited conceptual insight, while prototype-based approaches offer example-based explanations but often rely on rigid region selection and lack semantic consistency.
+To address these limitations, we propose PCMNet, a part-prototypical concept mining network that learns human-comprehensible prototypes from meaningful image regions without additional supervision.
+By clustering these prototypes into concept groups and extracting concept activation vectors, PCMNet provides structured, concept-level explanations and enhances robustness to occlusion and challenging conditions, which are both critical for building reliable and aligned AI systems.
+Experiments across multiple image classification benchmarks show that PCMNet outperforms state-of-the-art methods in interpretability, stability, and robustness.
+This work contributes to AI alignment by enhancing transparency, controllability, and trustworthiness in AI systems.
+Our code is available at: https://github.com/alehdaghi/PCMNet.
+Introduction
+Deep learning (DL) has advanced computer vision, enabling models to achieve remarkable performance across a wide range of tasks, including object detection and image classification (chai2021deep).
+However, despite these successes, DL models often operate as black boxes, and their lack of transparency has raised critical concerns regarding interpretability and reliability.
+This problem becomes even more significant in areas such as healthcare (fellous2019explainable), self-driving cars (kim2022xai), and security systems (ProtoPnet; PIPNet), where accurate predictions are not sufficient; it is crucial to understand the reasoning behind them.
+In these situations, explaining what the model is doing is important.
+It helps people trust the system, find problems if something goes wrong, and make sure the system follows important rules and laws (markus2021role).
+Therefore, researchers are now attempting to develop DL models that not only perform well but also provide interpretable and human-understandable explanations for their decisions (saranya2023systematic; vilone2020explainable).
+Common post-hoc explainability methods, such as GradCAM (chakraborty2022generalizing) and ScoreCAM (wang2020score), aim to explain a model’s decision by identifying which parts of the input image had the most influence on the output.
+These methods analyze the trained model’s gradients or activations to identify salient regions of the input.
+An example is shown in Fig.
+1 (a).
+Unlike these reverse-engineering methods, ante-hoc methods are designed to make the model explainable from the beginning.
+They guide the model to produce either global (he2025v2c; pcx; jiang2025knowledge) or local (PIPNet; ayoobi2025protoargnet; ProtoTree; zhu2024enhancing) interpretable parts that help explain its decisions.
+This idea is inspired by the recognition-by-components theory (biederman1987recognition), which suggests that humans understand objects by breaking them down into meaningful parts or concepts.
+Some methods, such as ProtoPNet (ProtoPnet) and PIPNet (PIPNet), as illustrated in Fig.
+1 (b), divide the spatial deep features (prior to the final pooling layer) into fixed, small patches and learn a set of prototypes from these regions.
+These prototypes, activated by specific input patches, serve as interpretable visual features that support the model’s decision.
+This also enables users to compare similar semantic concepts across different images, enhancing transparency and trust in the model’s reasoning capacity.
+However, these methods have notable limitations.
+Fixed patch sizes result in unstable prototype activation, particularly for larger semantic regions, where the patches fail to capture coherent or meaningful visual concepts.
+Conversely, very small patches lack sufficient contextual information, reducing their ability to represent interpretable or semantically rich components.
+Furthermore, discriminative and interpretable visual concepts often require larger regions to be represented clearly.
+When the model uses only small, fixed patches, it may fail to capture the full meaning of these concepts.
+This can lead to repetitive or shallow explanations, as seen in the bottom two rows of Fig.
+1 (b), where similar patterns appear multiple times.
+To address this, the regions used for prototypes should not be fixed.
+Instead, they should be adjustable and learned during training, allowing the model to focus on more meaningful and diverse parts of the image.
+Figure 1: Explanation methods.
 (a) Heatmap-based methods see the feature backbones as black-box models, and try to locate regions activated by the most important features.
-(b) With patch-based methods, the features are decomposed into components extracted by image patches, and an explanation is provided by these components.
-Direct application of naïve unsupervised part-discovery or Slot Attention-based methods often falls short in capturing primitive, semantically meaningful concepts from images and fails to provide intuitive explanations of model decisions based on extracted prototypes (van der Klis et al., 2023; Locatello et al., 2020).
-This limitation arises because part features are typically employed merely to identify similar parts across objects, without enforcing conceptual consistency, while Slot Attention features tend to lack explicit prototypical representations required for interpretable reasoning.
-The result, therefore, is a reduction in the diversity and a lack of interpretability of the prototypes, hindering the model’s ability to provide meaningful and comprehensive explanations for its decisions.
-To address these limitations, this paper introduces an intrinsically interpretable image classifier that mines meaningful and interpretable prototypical concepts from dynamic regions in images to explain the model’s decisions.
-Our unsupervised concept mining network identifies semantically meaningful components, relying only on class labels without requiring additional part labels or annotations.
-Our part-prototype concept mining network (PCMNet) detects larger, coherent regions of pixels as meaningful prototype parts.
-This design allows the components to identify relationships between semantically similar prototypes across instances from different classes.
-Then, we leverage these prototypes to mine a set of primitive yet discriminative concepts used to explain the model’s reasoning directly.
-For instance, Figure 2 demonstrates an use case where the prototype heatmaps, derived from our pipeline, shows that the decision-making was based on the ”lofty fenders”, ”shiny metal grille” and ”square shaped lights”, to conclude that the vehicle in the image, is a Dodge Durango.
-In the first stage, PCMNet employs an unsupervised part discovery mechanism to learn a set of prototypes using a novel center-clustering loss.
-In the second stage, the model encodes the deep features into a sparse Concept Activation Vector (CAV) (Kim et al., 2018), which provides a spatially and semantically interpretable explanation of the final decision.
+(b) Patch-based methods decompose the input into image patches and explain the decision using patch-derived components.
+(c) PCMNet, in contrast, mines both prototypical and non-prototypical concepts, offering a broader and more interpretable set of regions to explain the model decision.
+One possible direction is to discover meaningful regions using unsupervised part-based or slot-based mechanisms.
+However, direct application of naïve part-discovery methods (van2023pdisconet) or Slot Attention techniques (monet2021object; locatello2020slot) fails to produce semantically meaningful and interpretable components.
+These approaches typically aim to group similar visual parts across instances, but do not enforce conceptual consistency, resulting in ambiguous or fragmented representations (ProtoShare).
+Moreover, Slot Attention-based features often lack explicit prototypical structures, making it difficult to associate them with clear and intuitive explanations (li2021scouter).
+As a result, the learned prototypes tend to lack diversity and interpretability, limiting the model ability to generate meaningful justifications for its predictions.
+To address these limitations, we propose the Part-Prototypical Concept Mining Network (PCMNet), an interpretable image classifier that explains its predictions by extracting meaningful, part-based prototypical concepts  [^1] We define part-prototypes as localized visual regions that commonly appear across inputs (e.g., a car’s light), while the term concept describes a specific attribute or variation of that part (e.g., rectangular or circular shapes of the light).
+from selected regions.
+PCMNet first identifies relationships between semantically related prototypes across different instances, even across class boundaries and then derives a set of primitive class-discriminative concepts to form the basis for the model decision-making process.
+In the first stage, PCMNet uses an unsupervised part-discovery module that learns prototypes through a novel center-clustering loss.
+In the second stage, deep features are encoded into a sparse Concept Activation Vector (CAV) (CAV), which provides a spatially and semantically interpretable representation of the model’s decision.
 Each concept activation is computed as the cosine similarity between part features and class-aware prototype centroids.
-To ensure that these activated concepts are both primitive and class-discriminative, PCMNet applies a clustering algorithm to the part features corresponding to each class-specific prototype.
-The center of each resulting cluster then serves as a class-aware prototype centroid.
-Interpretation of the resulting CAV is straightforward, as it is produced by a sparse, non-negative linear layer that guarantees all concept values remain non-negative.
-The contributions of this paper are summarized as follows.
-(1) PCMNet is presented as a part-prototypical concept mining framework for explaining vision classifiers with primitive, prototypical, and interpretable human-meaningful components.
-(2) Concepts extracted by PCMNet are leveraged from a broader region than single patches to contain more semantical part-prototype abstraction, similar to the human brain, by using a novel contrastive prototype extraction.
-(3) Two-level clustering mechanisms are proposed for finding interpretable parts at the pixel level and primitive and intuitive concepts at the feature level.
-(4) Our experiments show that PCMNet explanation is more robust to occlusions, outperforming state-of-the-art patched-based prototype methods according to classification and explainability metrics.
-2. Related Work
-(a) Post-hoc or heatmap-based explanation: Heatmap-based explainability methods are a widely used family of post-hoc techniques that visualize which parts of an input image contribute most to a model’s decision.
+To ensure both conceptual simplicity and class-specific relevance, PCMNet applies a clustering algorithm to the part features associated with each prototype, and abstracts each cluster into a representative concept vector.
+The Concept Activation Vector (CAV) produced by PCMNet is inherently interpretable, as it is generated through a sparse, non-negative linear transformation that ensures all concept scores are positive.
+As illustrated in Fig.
+1 (c), PCMNet activates concepts such as “lofty fenders,” “shiny metal grille,” and “square-shaped lights” to support its prediction that the vehicle is a Dodge Durango, offering a clear and meaningful explanation of the model’s reasoning.
+[…]
+Related Work
+(a) Post-hoc or heatmap-based explanation.
+Heatmap-based explainability methods are a widely used family of post-hoc techniques that visualize which parts of an input image contribute most to a model’s decision.
 These techniques, often referred to as attribution methods, assign importance scores to different image regions to highlight influential features.
-Gradient-based methods, such as GradCAM (Selvaraju et al., 2017b) and ScoreCAM (Wang et al., 2020), generate heatmaps by backpropagating gradients concerning input features, revealing class-specific activation regions.
-While GradCAM produces class-dependent heatmaps, other methods like FullGrad (Srinivas and Fleuret, 2019) are class-agnostic, providing a more general view of model behavior across different outputs.
+Gradient-based methods, such as GradCAM (selvaraju2017grad) and ScoreCAM (wang2020score), generate heatmaps by backpropagating gradients concerning input features, revealing class-specific activation regions.
+While GradCAM produces class-dependent heatmaps, other methods like FullGrad (srinivas2019fullgrad) are class-agnostic, providing a more general view of model behavior across different outputs.
 Despite their popularity, gradient-based methods suffer from high sensitivity to noise, which can lead to unreliable and inconsistent heatmaps.
@@
 Beyond gradient-based methods, attribution propagation approaches offer an alternative way to decompose model predictions into layer-wise relevance scores.
-Techniques such as Layer-wise Relevance Propagation (LRP) (Bach et al., 2015) recursively distribute relevance through the network, providing a structured breakdown of model decisions.
-While initially designed for convolutional neural networks (CNNs), recent adaptations have extended these methods to vision transformers (ViTs) (Chefer et al., 2021), leveraging their self-attention mechanisms for improved interpretability.
-However, despite their effectiveness in highlighting important image regions, both gradient-based and attribution propagation methods remain fundamentally limited in providing high-level, human-interpretable concepts.
-Unlike our proposed PCMNet, which explains decisions through part-based prototypes, heatmap-based methods do not inherently capture semantic relationships between features, making them less suitable for interpretable concept-based explanations.
-Additionally, since these methods are inherently post-hoc, they do not enable real-time user intervention during test-time inference.
-This means that once a model makes a prediction, users cannot modify or refine the explanations interactively, limiting their practical use in applications requiring human-in-the-loop decision-making.
-In contrast, PCMNet introduces an ante-hoc, concept-driven interpretability approach that allows for more direct, structured, and interactive explanations by providing semantically meaningful components rather than relying solely on heatmap visualizations.
-(b) Slot attention: Slot Attention has emerged as a powerful mechanism for learning object-centric representations by dynamically assigning feature regions to a fixed number of latent slots (Locatello et al., 2020).
-Unlike traditional attention mechanisms that distribute importance weights across all features, Slot Attention iteratively refines a small set of slots, each representing a distinct object or meaningful component of an image.
-This allows models to decompose scenes into structured representations without explicit supervision (Kipf et al., 2022).
-By leveraging iterative updates with attention-based routing, Slot Attention enables more interpretable feature learning, where each slot encapsulates a semantically meaningful part of the input.
-This property makes it particularly useful for explainable AI, as it provides a structured decomposition of an image, rather than relying on uninterpretable distributed feature maps (Greff et al., 2020).
-Recent works have applied Slot Attention in various domains, such as object discovery, segmentation, and interpretable classification (Singh et al., 2022; Monet et al., 2021).
-In the context of explainable deep learning, it has been used to learn disentangled representations where each slot corresponds to a distinct semantic concept (Karazija et al., 2021).
-However, these methods typically focus on object-level decomposition rather than part-based representation.
-Our PCMNet builds on this idea by using Slot Attention-inspired mechanisms to discover meaningful part prototypes rather than full objects, ensuring better granularity in concept-based explanations.
-By applying slot-like iterative attention within a contrastive prototype extraction framework, PCMNet refines part-based representations, making them more interpretable and aligned with human perception and reasoning.
-(c) Concept-based explanation: Contrary to traditional local feature attribution methods that assess the importance of input-level features, concept-based XAI methods (Dreyer et al., 2024; Chen et al., 2019; Rymarczyk et al., 2021; Bach et al., 2015) analyze the role of latent representations in specific layers of a deep neural network (DNN).
-These methods explore concepts through various mechanisms, including individual neurons, directional representations such as Concept Activation Vectors (CAVs) (Kim et al., 2018), or broader feature subspaces.
-Early XAI research primarily focused on understanding how these concepts contribute to global decision-making, identifying the most relevant concepts for a given output class (Dreyer et al., 2024; Bach et al., 2015).
-Some approaches enhance model interpretability by employing local feature attribution techniques to localize and quantify the importance of concepts for individual predictions, thereby enabling concept-based explanations at the instance level (Nauta et al., 2023, 2021).
+Techniques such as Layer-wise Relevance Propagation (LRP) (bach2015lrp) or KD-FMV (jiang2025knowledge) recursively distribute relevance through the network, providing a structured breakdown of model decisions.
+While initially designed for convolutional neural networks (CNNs), recent adaptations have extended these methods to vision transformers (ViTs) (chefer2021transformer), using their self-attention mechanisms for improved interpretability.
+However, despite their effectiveness in highlighting important image regions, both gradient-based and attribution propagation methods remain fundamentally limited in providing high-level, human interpretable concepts.
+Unlike our PCMNet, which explains decisions through part-based prototypes, heatmap-based methods do not inherently capture semantic relationships between features, making them less suitable for interpretable concept-based explanations.
+(b) Concept-based explanation.
+Concept-based XAI methods (pcx; ProtoPnet; ProtoShare; bach2015lrp; he2025v2c) analyze the role of latent representations in specific layers of a deep neural network to explore concepts from input images and find Concept Activation Vectors (CAVs) (CAV) for the made decision.
+Early XAI research primarily focused on understanding how these concepts contribute to global decision-making, identifying the most relevant concepts for a given output class (pcx; bach2015lrp).
+Some approaches enhance model interpretability by employing local feature attribution techniques to localize and quantify the importance of concepts for individual predictions, thereby enabling concept-based explanations at the instance level (PIPNet; ProtoTree; ayoobi2025protoargnet).
 While these methods provide faithful explanations aligned with the model’s internal reasoning, they are vulnerable to performance degradation under occlusion, as their extracted concepts often rely on limited and highly localized discriminative regions of the input.
-PCMNet addresses this limitation by part-based prototypical concept reasoning to rely on other discriminative parts that are not occluded.
-Additionally, although instance-wise concept-based explanations provide deeper insights, they can also be overwhelming and difficult for stakeholders to process, as hundreds of concepts may need to be analyzed for each instance.
-To address this, some works highlight the benefits of visualizing local decisions within a global embedding space (Dreyer et al., 2024; Donnelly et al., 2022).
-With PCMNet, we build upon this idea by introducing part prototypes that encapsulate local concepts.
-This approach reduces the cognitive load for human interpretation by enabling comparisons between individual (local) prediction and prototypical concepts, making explanations more intuitive and structured.
-3. Proposed Prototypical Concept Mining Network
-Figure 3.
-Overall architecture of PCMNet.
+PCMNet addresses this limitation by wider range of part-based prototypical concept reasoning to rely on other discriminative parts that are not occluded.
+Recently, Concept Bottleneck Models (CBMs) has emerged as a research avenue for inherently interpretable approaches that predict intermediate human-understandable concepts before class prediction (xie2025discovering; posthocCBM; rao2024discover; he2025v2c).
+Unlike patch- or part-prototype models, CBMs explicitly separate the reasoning process into concept prediction and concept-to-label mapping, allowing direct human intervention on the concept layer.
+However, their reliance on annotated concepts limits scalability, and the fidelity of the bottleneck depends on concept quality.
+Recent methods, including post-hoc CBMs (he2025v2c; xie2025discovering) and vision-to-concept tokenizers that learn visual concept vocabularies without text supervision—aim to bridge this gap by enabling scalable and more interpretable reasoning.
+Prototypical Concept Mining Network
+(a) PCMNet
+(b) CMM
+Figure 2: (a) Overall architecture of PCMNet.
 In the first stage, part prototypes are learned from spatial features extracted by the backbone.
@@
 This approach enables interpretable and concept-driven classification.
-To explain the decision for a classification problem with a training set $\mathcal{T}$ containing $N$ images $\{(x^{1},y^{1}),...,(x^{N},y^{N}\}\in\mathcal{X}\times\mathcal{Y}$, with $L$ classes, the PCMNet extracts primitive and sparse concepts that are activated and discriminate the object.
-These concepts are described in a limited set of prototypes that existed for almost all images in the datasets (e.g., lights, fenders, and grilles for cars or heads and wings for birds).
-In other words, the output of our model is the activated concept vector $z_{i}=[z_{p}^{i}]_{p=1}^{K}$, in which $p$ is the prototype index.
+(b) Concept Mining Module.
+First, a part centroid list is generated by applying DBSCAN(DBSCAN) clustering within each class and computing the centers of the resulting clusters.
+These centroids serve as frequently occurring concept prototypes.
+Next, to activate the most relevant concept for a given part feature, the similarity (inverse distance) to the centroids is measured.
+These values are then used as CAVs to inform the model’s decision-making.
+To explain classification decisions over a training set $\mathcal{T}={(x^{1},y^{1}),...,(x^{N},y^{N})}\subset\mathcal{X}\times\mathcal{Y}$ with $L$ classes, PCMNet extracts sparse and primitive concepts that are both activated and discriminative for object recognition.
+These concepts are captured through a compact set of prototypes that recur across images in the dataset (e.g., lights, fenders, and grilles for cars; heads and wings for birds).
+The model outputs a class-based Concept Activation Vector (CAV) $\mathbf{z}^{i}=[\mathbf{z}^{i}_{1};\dots;\mathbf{z}^{i}_{K}]$, where ; denotes concatenation and each $\mathbf{z}_{p}$ represents the activated concepts from part $p$.
 Fig.
-3 shows the overall overview of our approach.
-PCMNet transforms features extracted from the backbone model into explainable and interpretable concepts in three steps.
-Step 1: The model finds semantically independent, meaningful regions from input images in an unsupervised way by minimizing our novel Marginal Cluster Center loss to learn the prototypes that describe these parts.
-Then, the part features are disentangled to discriminate the class label.
-Step 2: Once the part features are learned for the training data, for each part inside all instances of each class, a set of shared elementary characteristics is found by applying a clustering approach to these features.
-Then, the centroid list, which represents the prototypical concepts, is generated by calculating the center of each cluster.
-Step 3: Extract the Concept Activation Vector (CAV) for the generated centroid list by computing the similarity of extracted features and each centroid.
-The CAVs are used to discriminate the class of the input and explain the decision.
-Also, to avoid losing some discriminative concepts that are not prototypical (i.e., specific and unique for this class and not available for others), we combine the features that are located outside of prototypical maps into the CAVs for classification.
-3.1. Step 1: Part-Prototype Discovery
+2 shows the overall overview of PCMNet, which transforms features extracted from the backbone into explainable and interpretable concepts in three steps.
+Step 1: PCMNet identifies semantically meaningful and independent regions from input images in an unsupervised manner by minimizing the Marginal Cluster Center loss ($\mathcal{L}_{\text{mcc}}$).
+This process learns part-prototype features $\mathbf{f}^{i}_{p}$ that are informative for class discrimination.
+Step 2: Once part-level features are obtained, the Concept Mining Module (CMM) $\mathcal{H}$ aggregates these features across all instances of each class to discover a set of shared elementary visual patterns.
+Cluster centers derived from this process form the list $\mathcal{C}$, representing the core prototypical concepts.
+Step 3: For each input, a Concept Activation Vector (CAV) $\mathbf{z}^{i}$ is computed by measuring the similarity between its extracted part features and the concept list $\mathcal{C}$.
+These CAVs are used both for classification and for explaining the model’s decision.
+To preserve class-specific but non-shared discriminative cues, PCMNet also incorporates features outside the prototypical regions into the CAVs.
+Step 1: Part-Prototype Discovery
 Our model is shown in Fig.
-3.
-The input images are given to a backbone to extract the 3D deep features as $F^{i}\in\mathbb{R}^{D\times w\times h}$ where $D$ is the channel size and $w,h$ denotes the width and height.
+2.
+The input images are given to a backbone to extract the 3D deep features as $F^{i}\in\mathbb{R}^{D\times w\times h}$ where $D$ is the channel size and $w,h$ denote the width and height.
 Based on extracted deep features $F^{i}$, our model finds semantically independent meaningful regions that can be shared in different inputs.
@@
 For determining regions, the model scores each pixel in $F^{i}$ to specify their belongingness to each part-prototype class.
-The mapping score for each pixel in the location $(q)$ is represented by $M^{i}$, we have $\sum_{p=1}^{K+1}M^{i}_{p}(q)=1$.
+The mapping score for each pixel $(a)$ is represented by $M_{p}^{i}(a)$, we have $\sum_{p=1}^{K+1}M^{i}_{p}(a)=1$.
 The spatial feature for each part, $F^{i}_{p}$, is computed by element-wise multiplication of $M^{i}_{p}$ to $F^{i}$.
-The part features, $\mathbf{f}^{i}_{p}$ is computed as:
-(1) $\mathbf{f}^{i}_{p}=\mathcal{GP}(\texttt{conv}(F^{i}_{p}))\in\mathbb{R}^{d_{f}},$
+Then part features, $\mathbf{f}^{i}_{p}$ is computed as:
+$\vskip-2.84544pt\mathbf{f}^{i}_{p}=\mathcal{GP}(\texttt{conv}(F^{i}_{p}))\in\mathbb{R}^{d_{f}},$ (1)
 where $\mathcal{GP}$ is the global average pooling and $d_{f}$ is the dimension of features.
-The last index (K+1) is used to cover the background or non-part regions of the foreground (Alehdaghi et al., 2025).
+The last index (K+1) is used to cover the background or non-part regions of the foreground (BMDG).
 We name these features non-prototypical concepts that are computed as:
-(2) $\mathbf{g}^{i}=\mathcal{MP}(\texttt{conv}(F^{i}_{K+1}))\in\mathbb{R}^{d_{f}},$
+$\vskip-5.69046pt\mathbf{g}^{i}=\mathcal{MP}(\texttt{conv}(F^{i}_{K+1}))\in\mathbb{R}^{d_{f}},$ (2)
 where $\mathcal{MP}$ is the max-pooling.
-We use the PDiscoNet (van der Klis et al., 2023) as the baseline for extracting parts features.
-Our part loss comprises their final loss ($\mathcal{L}_{\text{bl}}$) without their orthogonal component and our marginal cluster center loss ($\mathcal{L}_{\text{mcc}}$):
-(3) $\mathcal{L}_{\text{part}}=\mathcal{L}_{\text{bl}}+\alpha\mathcal{L}_{\text{mcc}}.$
-The goal of $\mathcal{L}_{\text{mcc}}$ is to ensure the selected region for each part remains consistent across all inputs and serves as a prototype; it must capture similar semantic information across different classes.
-Additionally, these sub-part features should be both class-descriptive and contain unique ID-aware information.
-3.1.1. Marginal Cluster Center Loss
-We proposed learnable prototypes as features to be distant from each other and then pushed the model to discover the parts from input images to be close to these prototypes.
-A soft margin is used to make space for the same part from input images with different classes.
-(4) $\mathcal{L}_{\text{mcc}}=\sum_{p=1}^{K}[||\mathbf{f}^{i}_{p}-c_{p}||-m_{1}]_{\star}+\frac{1}{K}\sum_{q=p+1}^{K}[m_{2}-||c_{p}-c_{q}||]_{\star},$
-where $c_{p}$ is a learnable prototype for part $p$ and $[\cdot]_{\star}$ is $\max(\cdot,0)$.
-3.2. Step 2: Part-Centroid List Generation
-During training in Step 1, the backbone learns to extract and use part-prototype features from the input data to fulfill its training task.
-To find all relevant spatial and class-aware concepts, as shown at the top of Fig.
-4, we collect all extracted features from the backbone for each class and each part and then cluster them to have a minimal set of plain concepts that consistently describe the visual attributes of those parts.
-Multiple concepts are considered.
-They can be seen in each part of each class for more complex objects.
-To mine those concept prototypes, the centroids of each cluster are computed as $C(l,p,j)\in\mathbb{R}^{d}_{f}$ for a clustering index of $l$, part $p$, and a class of $j$.
-3.3. Step 3: Concept Activation Mining
+We use the PDiscoNet (van2023pdisconet) model for extracting parts features.
+Our part loss comprises their final part-discovery loss ($\mathcal{L}_{\text{pd}}$) without their orthogonal component and with our marginal cluster center loss ($\mathcal{L}_{\text{mcc}}$):
+$\vskip-4.26773pt\mathcal{L}_{\text{part}}=\mathcal{L}_{\text{bl}}+\alpha\mathcal{L}_{\text{mcc}}.$ (3)
+where $\alpha$ is a parameter regularizing the effect of $\mathcal{L}_{\text{mcc}}$.
+The goal of $\mathcal{L}_{\text{mcc}}$ is to ensure that each selected region is consistent across inputs and captures semantically similar content, while remaining class-discriminative and identity-aware.
+Marginal Cluster Center Loss
+To ensure part-prototypes are both distinct and semantically meaningful, we encourage separation in the feature space while guiding part-level features to align with their corresponding prototypes.
+To handle intra-part variation across classes, we apply a soft-margin loss that allows features from the same part (but different classes) to stay within a bounded distance of their prototype.
+This encourages the model to discover consistent, class-aware part structures while preserving inter-class separation between prototypes.
+The loss can be written as
+$\mathcal{L}_{\text{mcc}}\!=\!\sum_{p=1}^{K}[|\mathbf{f}^{i}_{p}-c_{p}|\!-\!m_{1}]_{\star}\!+\!\frac{1}{K}\!\sum_{q\neq p}^{K}[m_{2}\!-\!|c_{p}-c_{q}|]_{\star},$ (4)
+where $c_{p}$ is a learnable prototype and $\cdot$$_{\star}$ is $\max(\cdot,0)$.
+Step 2: Part-Centroid List Generation
+In Step 1, the backbone learns to extract meaningful part-level features that support the classification task.
+To discover compact and consistent class-specific concepts, we collect all part features from the training set.
+Specifically, for each class $j$ and part $p$, we construct $B^{j}_{p}=\{\mathbf{f}_{p}^{i}\mid y^{i}=j\}\in\mathbb{R}^{N_{j}\times d_{f}}$, where $N_{j}$ is the number of training samples from class $j$.
+To identify semantically coherent patterns, we apply the clustering algorithm separately to each $B^{j}_{p}$, yielding $T^{j}_{p}\in\mathbb{N}^{N_{j}}$, which assigns cluster labels to the part features.
+We then compute the centroid of each cluster $l$ (representing a visual concept) as:
+$\vskip-2.84544pt\mathcal{C}(l,p,j)=\text{mean}(\{B^{j}_{p}[T^{j}_{p}=l]\})\in\mathbb{R}^{d_{f}}.$ (5)
+The total number of discovered concepts across all classes and parts is computed as $d_{c}=\sum_{j=1}^{L}\sum_{p=1}^{K}\max(T^{j}_{p})$.
+This process, shown in the top of Fig.
+2(b), allows the model to build a compact and expressive vocabulary of class-specific, part-aware visual concepts.
+Step 3: Concept Activation Mining
 Once the centroid list is generated, each concept’s activation value is defined by the similarity of extracted features to each centroid:
-(5) $\mathbf{z}^{i}=[\mathcal{S}(C(l,p,j),\mathbf{f}_{p}^{i})\;\forall p\in\{1..K\}\;\text{and}\;j\in\{1..L\}]\in\mathbb{R}^{d_{c}}$
-where $\mathcal{S}$ is cosine distance and [.][.][ .
-] represents the vector concatenation.
-$d_{c}$ is the number of total concept centroids.
+$\mathbf{z}^{i}=\{\mathcal{S}(\mathcal{C}(l,p,j),\mathbf{f}_{p}^{i})\;\forall p\in\{1..K\}\;\text{and}\;j\in\{1..L\}\},$ (6)
+where $\mathbf{z}^{i}\in\mathbb{R}^{d_{c}}$ ans $\mathcal{S}$ is cosine distance.
 We aim to explain each output class of our model using a small set of interpretable concepts.
@@
 Here, $W_{1}\in\mathbb{R}^{d_{c}\times L}$, $W_{2}\in\mathbb{R}^{d_{f}\times L}$ and $b\in\mathbb{R}^{L}$ denote the classification weights and bias term, respectively.
-This layer is trained with the following classification loss, using the GLM-SAGA optimizer (Wong et al., 2021):
-(6) $\mathcal{L}_{\text{cl}}=\mathcal{L}_{ce}(W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b,y^{i})+\lambda\mathcal{R}(W_{1}),$
-where $\mathcal{L}_{\text{ce}}$ is the cross-entropy loss, $y^{i}$ is the label, $\lambda$ is the sparsity regularization strength and $\mathcal{R}(W)=(1-\gamma)\frac{1}{2}\|W\|_{F}+\gamma\|W\|_{1}$, here $\|W\|_{F}$ is the Forbenius norm and $\|W\|_{1}$ is the element-wise matrix.
-Figure 4.
-Concept Mining Module.
-First, a part centroid list is generated by applying DBSCAN(Ester et al., 1996) clustering within each class and computing the centers of the resulting clusters.
-These centroids serve as frequently occurring concept prototypes.
-Next, to activate the most relevant concept for a given part feature, the similarity (inverse distance) to the centroids is measured.
-These values are then used as CAVs to inform the model’s decision-making process .
-3.4. Overall Training
-We jointly optimize the network in an end-to-end manner by loss:
-(7) $\mathcal{L}=\mathcal{L}_{\text{part}}+\beta\mathcal{L}_{\text{cl}},$
-where $\beta$ is hyperparameter. For the first step of training, we set $\beta=0$; then, once the part-prototypical concepts are learned, we set $\beta=2$ based on our hyperparameter analysis.
-4. Experiments and Results
+This sparse layer is trained with the following sparse classification loss (wong2021leveraging):
+$\mathcal{L}_{\text{cl}}=\mathcal{L}_{ce}(W_{1}^{T}\mathbf{z}^{i}+W_{2}^{T}\mathbf{g}^{i}+b,y^{i})+\lambda\mathcal{R}(W_{1}),$ (7)
+where $\mathcal{L}_{\text{ce}}$ is the cross-entropy loss, $y^{i}$ is the label, $\lambda$ is the sparsity regularization strength and $\mathcal{R}(W)=(1-\gamma)\frac{1}{2}\|W\|_{F}+\gamma\|W\|_{1}$, here $\|W\|_{F}$ is the Forbenius norm and $\|W\|_{1}$ denotes the element-wise matrix.
+Overall Training
+We jointly optimize the network in an end-to-end manner:
+$\mathcal{L}=\mathcal{L}_{\text{part}}+\beta\mathcal{L}_{\text{cl}},$ (8)
+where $\beta$ is hyperparameter. We set $\beta=0$ during initial training and $\beta=2$ after learning part-prototypical concepts.
+Experiments and Results
 Method Consistency (Intra) $\uparrow$ Consistency (Inter) $\downarrow$ F(1)-F(5) $\uparrow$ Sp $\uparrow$ Stability C Acc $\uparrow$
-Stanford Cars Resnet50+Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 0.16 - 0.23 22.74 65.4 84.1
-Resnet50+W $\times$ Act 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 1.54 - 6.61 22.74 65.4 84.1
-ProtoPNet + Act 45.70$\pm$4.9 11.22$\pm$5.5 2.2-55.64 60.92 69.6 84.5
-ProtoPNet + W $\times$ Act 45.70 $\pm$ 4.9 11.22$\pm$5.5 9.3-80.12 60.92 69.6 84.5
-PiPNet(Covx) + Act 52.78$\pm$7.4 9.07$\pm$9.4 0.02-60.62 70.59 65.2 88.5
-PiPNet(Covx) + W $\times$ Act 52.78 $\pm$ 7.4 9.07$\pm$9.4 10.83-95.66 70.59 65.2 88.5
-PiPNet(Res50) + Act 42.40$\pm$17.5 17.4$\pm$2.4 0.18 - 1.21 63.19 60.8 86.46
-PiPNet(Res50) + W $\times$ Act 42.40$\pm$17.5 17.4$\pm$2.4 9.43 - 93.15 63.19 60.8 86.46
-Ours + Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 37.94 - 63.3 57.45 71.5 88.7
-Ours + W $\times$ Act 55.32$\pm$6.1 1.88 $\pm$ 9.37 59.33 - 91.73 57.45 71.5 88.7
-Birds (CUB 200) Resnet50+Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 0.35 - 1.12 24.14 60.3 81.2
-Resnet50+Rel $\times$ Act 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 2.43 - 5.73 24.14 60.3 81.2
-ProtoPNet + Act 51.47$\pm$5.8 10.16$\pm$9.2 1.15-51.96 79.51 63.6 81.45
-ProtoPNet + W $\times$ Act 51.47 $\pm$ 5.8 10.16$\pm$9.2 9.36-75.50 79.51 63.6 81.45
-PiPNet(Covx) + Act 54.80$\pm$8.3 7.77$\pm$8.0 0.01-54.14 80.72 64.7 85.0
-PiPNet(Covx) + W $\times$ Act 54.80 $\pm$ 8.3 7.77$\pm$8.0 5.13-92.83 80.72 64.7 85.0
-PiPNet(Res50) + Act 42.28$\pm$3.2 15.57$\pm$15.7 0.14 - 2.35 77.83 65.9 82.0
-PiPNet(Res50) + W $\times$ Act 42.28$\pm$3.2 15.57$\pm$15.7 9.40 - 89.77 77.83 65.9 82.0
-PCMNet (Ours) + Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 41.82 - 75.24 57.45 67.1 84.7
-PCMNet (Ours) + W $\times$ Act 58.55$\pm$7.5 1.22 $\pm$ 10.1 55.13 - 89.88 57.45 67.8 84.7
-Table 1. Performance of PCMNet according to xAI metrics on the Stanford Cars and Birds datasets.
-PCMNet is validated on standard benchmarks in prototype literature: Stanford Cars (Krause et al., 2013) (196 classes of different make and models) and CUB-200-2011 (Wah et al., 2011) (200 bird species, including attribute annotation for visual features).
-The ResNet50 model pretrained on ImageNet (Deng et al., 2009) is used as the feature backbone.
-Each training batch (32 images) is resized to 488×488, randomly cropped, erased, and padded.
-Optimization follows ADAM (Kingma and Ba, 2014) with a linear warm-up.
-We set regularization parameters $\lambda$ and $\gamma$ as same in (Oikarinen et al., 2023).
-$\alpha=1.5$, $\beta=2$, with $m_{1}=0.3$ and $m_{2}=1.5$ chosen via hyperparameter analysis in the suppl.
-material.
-More details on the experimental protocol are presented in the supplementary materials.
-4.1. Explainability Metrics
-To evaluate the explainability of final features, we rely on performance metrics established in the literature (Dreyer et al., 2024) with some minor modifications to express better comprehension.
-(a) Faithfulness:: Faithfulness(Dasgupta et al., 2022; Chan et al., 2022) evaluates how much the activated concepts influence the model’s final prediction, serving as a key metric for assessing the quality of explanations.
+Cars-196
+Baseline 83.45 $\pm$ 2.7 27.26 $\pm$ 28.10 1.54 - 6.61 22.74 65.4 84.1
+[…]
+[…]
+[…]
+[…]
+[…]
+CUB-200
+Resnet50 57.38 $\pm$ 3.50 25.45 $\pm$ 23.82 2.43 - 5.73 24.14 60.3 81.2
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 1: Performance of PCMNet according to xAI metrics on the Stanford Cars and Birds datasets.
+Implementation Details
+PCMNet is evaluated on two standard prototype-learning benchmarks: Stanford Cars (scar) (196 car classes) and CUB-200-2011 (cubs) (200 bird species).
+We use a pretrained ResNet50 as the backbone.
+Each training batch (32 images) is resized to 488×488 with random cropping and padding.
+Optimization is performed using Adam (kingma2014adam).
+Regularization parameters $\lambda$ and $\gamma$ follow (oikarinen2023label); we set $\alpha=1.5$, $\beta=2$, with margins $m_{1}=0.3$ and $m_{2}=1.5$ based on hyperparameter tuning analysis [^2] Full paper: https://arxiv.org/pdf/2504.12197.
+End-to-End Multi-Stage Training: PCMNet modules (part discovery, concept clustering, and concept activation) are trained end-to-end in a unified pipeline over 40 epochs:
+- Step 1 (initial epochs): The part discovery module is optimized to localize semantically meaningful part features. This stage continues until convergence of $\mathcal{L}_{\text{part}}$.
+- Step 2 (intermediate epochs): Concept clustering runs intermittently (every 5 epochs) to reduce computation while keeping centroids up to date with changing representations.
+- Step 3 (final epochs): The Concept Activation Mining Module (CMM) is trained alongside the backbone, using the updated centroids from Stage 2 to align activated features with discriminative concept prototypes.
+At inference, PCMNet behaves as a standard forward-pass classifier.
+Explainability Metrics
+To evaluate explainability, we use established metrics from (pcx), with minor adjustments.
+(a) Faithfulness: Faithfulness (fait1; fait2) evaluates how much the activated concepts influence the model’s final prediction, serving as a key metric for assessing the quality of explanations.
 One widely used approach for measuring faithfulness is concept deletion, which involves removing selected concepts from the model’s reasoning process and observing the resulting change in output confidence.
-Prior works, such as (Dreyer et al., 2024), typically assess this by removing only the single most important concept.
+Prior works, such as (pcx), typically assess this by removing only the single most important concept.
 However, this limited scope fails to capture the broader contribution of other activated concepts and may not reveal the full extent of their impact.
@@
 (b) Stability: The extracted concepts from input images that share some semantic concepts must be stable and explainable for unseen images.
-To evaluate stability, we compute CAVs on $k$-fold subsets of the data ($k$ = 10 as default) similar to (Dreyer et al., 2024) for all classes at steps 2 and 3.
+To evaluate stability, we compute CAVs on $k$-fold subsets of the data ($k$ = 10 as default) similar to (pcx) for all classes at steps 2 and 3.
 We then map CAVs together using a Hungarian loss function and measure the cosine similarity between them.
@@
 This is because a uniform distribution of concept activations would provide little information on the importance of specific concepts or parts of the images, i.e., a high entropy in the generated concepts.
-4.2. Comparison to Prototype xAI Methods
-To evaluate PCMNet, we report the classification accuracy and the XAI evaluation criteria discussed in Sec 4 and compare them to other ante-hoc methods ( ProtoPNet(Chen et al., 2019) and PIPNet(Nauta et al., 2023)) in tables 1.
-The consistency metric is reported in two ways: Intra and Inter, which are cosine similarities between activated concepts of the same class and different classes, respectively.
+Comparison to Prototype xAI Methods
+To evaluate PCMNet, we report classification accuracy and XAI metrics, and compare them with other ante-hoc methods (ProtoPNet(ProtoPnet), PIPNet(PIPNet), ProtoViT (protoViT) and MCPNet (Wang2024MCPNet)) in Table 1.
+The consistency metric is reported in two forms: Intra and Inter, measuring the cosine similarity between activated concepts within the same class and across different classes, respectively.
 Faithfulness is reported as F(n), measuring accuracy drops when the top n important concepts are removed from the decision.
-To compute the contribution of each concept or feature to the final decision, we evaluate two strategies: (1) using raw values of each concept (”Act”), and (2) computing a weighted contribution ( gradient) as the product of raw values and its corresponding weight in the classification layer (“W $\times$ Act”).
-4.3. Ablation Studies
-4.3.1. Number of Prototypes
-We assess the impact of the number of parts on both classification accuracy and faithfulness in Table 2.
-As the model extracts more prototypical concepts, its decisions become more aligned with the underlying features, enhancing interpretability.
-Interestingly, when the number of prototypes is smaller, the faithfulness scores for the top concepts tend to be higher.
-This is because the model relies on a more compact set of prototypes, concentrating its decisions on fewer, more influential concepts.
-K value Cars Birds
-C Acc F(1)-F(5) C Acc F(1)-F(5)
-0 (Resnet50) 84.1 1.54 - 6.61 83.2 1.21-5.56
-4 83.5 79.96 - 91.01 83.5 71.42-85.64
-5 84.1 70.11 – 91.33 84.9 66.7 - 85.9
-6 85.2 66.02 – 90.65 85.7 59.9 - 86.7
-7 86.7 62.47 – 91.25 86.0 55.7 - 87.3
-8 87.5 59.33 – 91.73 85.1 53.0 - 86.4
-9 86.4 55.10 – 91.52 84.4 52.5-86.1
-Table 2. Impact on performance of the number of prototypes.
-Figure 5. Visualizing activated concepts in the test set and examples from the training set for Stanford Cars and CUB datasets.
-4.3.2. Effect of Modules
-To evaluate the effectiveness of each component in PCMNet, we conduct an ablation study on the baseline PDiscoNet model, which extracts local parts without explicit concept learning.
-As shown in Table 3, we analyze classification accuracy (C Acc) and feature faithfulness under occlusion (F(3)) on both the Cars and Birds datasets, alongside the model size and FLOPs.
-Introducing the Marginal Clustering Center (MCC) loss into the baseline improves classification performance (from 84.2% to 86.1% on Cars) and more than doubles faithfulness scores, indicating that enforcing clustering consistency across part features supports more explainability.
-The inclusion of the Concept Mining Module (CMM)—responsible for extracting sparse and interpretable concepts—dramatically boosts faithfulness (from 4.9% to 55.6% on Birds), even with minimal impact on accuracy.
-When both MCC and CMM are combined, PCMNet achieves the best results across all metrics, including 87.5% classification accuracy and 67.4% faithfulness on Cars, with only a marginal increase in parameter count and computational cost.
-These findings underscore the complementary roles of MCC in enforcing cluster-level consistency and CMM in mining high-quality, human-aligned concepts.
+To compute the contribution of each concept to the final decision, we compute a weighted score as the product of its raw value and its corresponding classification weight.
+Ablation Studies
+(a) Effect of Number of Concept Center centroids ($d_{c}$):
+The size of the concept list ($d_{c}$) determines the semantic resolution of PCMNet and is influenced by both the number of parts and the number of training classes.
+As the number of parts ($K$) increases, so does the total number of extracted centroids.
+We analyze this dependency further in Appendix B2.
+While richer concept sets can improve interpretability, they also increase memory and computational cost.
+To manage this, we apply a post-clustering refinement step that merges similar centroids within or across classes using hierarchical clustering (murtagh2011ward).
+Table 2 reports the effect of varying the merging level and threshold—defined as a percentage of the maximum inter-centroid distance—on the Cars dataset.
+We find that moderate merging (e.g., Level 1 with 10% threshold) significantly reduces $d_{c}$ with minimal impact on classification accuracy or faithfulness (F(3)).
+However, excessive merging (e.g., Level 3 at 10%) leads to performance degradation, suggesting the importance of controlled prototype compression.
+Threshold (%) Level $d_{c}$ C Acc (%) F(3) (%)
+0% 1 3092 90.2 67.4
+2 2473 85.7 65.7
+3 698 84.4 66.5
+5% 1 2862 89.5 66.5
+2 1805 83.6 66.2
+3 645 80.5 65.7
+10% 1 1657 86.2 67.9
+2 1263 82.7 66.9
+3 454 74.7 65.1
+Table 2: Impact of hierarchical clustering thresholds and levels on the number of concept prototypes ($d_{c}$), classification accuracy, and interpretability (faithfulness F(3)) on the Cars dataset.
+Thresholds, expressed as a percentage of the maximum pairwise centroid distance.
+(b) Effect of Modules:
+To assess the impact of each PCMNet component, we conduct an ablation study on the baseline PDiscoNet model without concept learning.
+Additionally, the effect of using different types of concept vectors—prototypical and non-prototypical—is examined in Appendix B2.
+Table 3 reports classification accuracy (Acc) and faithfulness (F(3)) on the Cars and Birds datasets.
+Adding the Marginal Clustering Center (MCC) loss improves accuracy (e.g., from 81.2% to 87.8% on Cars) and doubles faithfulness, highlighting the benefits of enforcing cluster-level consistency.
+Incorporating the Concept Mining Module (CMM) yields a major boost in faithfulness (from 4.9% to 55.3% on Birds), with minimal cost to accuracy.
+Combining both modules yields the best overall performance: 90.2% accuracy and 67.4% faithfulness on Cars, with only +0.4M parameters and +0.3G FLOPs.
+Compared to recent interpretable baselines like PipPNet, ProtoPNet, ProtoViT, and MCPNet, PCMNet achieves a better balance between accuracy, faithfulness, and efficiency, especially outperforming ProtoViT in F(3) (+22.8%) and requiring fewer FLOPs (17.4G vs.
+24.8G).
+This confirms the complementary value of MCC and CMM in delivering interpretable and efficient visual recognition.
+The trade-off information is discussed in the ArXiv version.
 Settings Cars Birds #Params FLOPs
-C Acc F(3) C Acc F(3)
-Baseline(van der Klis et al., 2023) 81.2 5.4 82.3 4.9 27.1M 17.2G
-Base + MCC 86.1 9.6 84.5 9.1 27.2M 17.2G
-Base + CMM 84.3 59.2 82.1 55.6 27.4M 17.5G
-Base + MCC + CMM 87.5 67.4 84.7 64.9 27.5M 17.5G
-PipPNet(Resnet) 86.4 62.7 82.0 58.6 25.1M 27.6G
-PipPNet(Covx) 88.5 60.6 85.0 57.5 29.3M 32.3G
+Acc F(3) Acc F(3)
+Baseline 81.2 5.4 82.3 4.9 27.1M 17.1G
+Baseline + MCC 87.8 9.2 84.9 8.9 27.2M 17.1G
+Baseline + CMM 85.8 58.6 82.5 55.3 27.4M 17.4G
+PCMNet (Full) 90.2 67.4 85.1 64.9 27.5M 17.4G
+PiPNet 86.4 62.7 82.0 58.6 25.1M 27.6G
 ProtoPNet 84.5 38.6 81.4 31.7 28.9M 18.4G
-Table 3. Classification accuracy and faithfulness using different PCMNet modules.
-Figure 6. t-SNE visualization of concept-level part features extracted by PCMNet. Colors represent different prototypes, while shapes indicate object classes.
-4.4. Robustness to Occlusion
-To evaluate the robustness of PCMNet’s decision-making under occlusion, we conducted controlled experiments in which regions corresponding to the most activated concepts were systematically masked, and the resulting images were re-evaluated by the model.
+ProtoViT 91.8 44.6 85.1 36.7 22.1M 24.8G
+MCPNet 80.1 64.0 80.1 39.3 24.6M 17.4G
+Table 3: Classification accuracy and faithfulness using different PCMNet modules.
+Robustness to Occlusion
+To evaluate PCMNet robustness under occlusion, we conducted controlled experiments by masking regions linked to the most activated concepts and re-evaluating the modified images with the model.
 This experiment aims to assess the stability of model predictions when key interpretable regions are removed.
@@
 We consider three levels of occlusion, masking 10%, 20%, and 30% of the image area.
-PCMNet’s performance under these conditions is compared against two other interpretable baselines: ProtoPNet (Chen et al., 2019) and PIPNet (Nauta et al., 2023).
-Figure 7. Classification Accuracy and Faithfulness under Different Occlusion Levels.
+PCMNet’s performance under these conditions is compared against two other interpretable baselines: ProtoPNet (ProtoPnet) and PIPNet (PIPNet).
+(a) Accuracy.
+(b) Faithfulness F(3)
+Figure 3: Classification accuracy and faithfulness for a growing level of occlusion.
 Fig.
-7 shows that PCMNet maintains stable accuracy at lower occlusion levels and degrades more gradually than ProtoPNet and PIPNet.
-This resilience stems from PCMNet’s broader part-based concepts, which introduce redundancy and enable accurate predictions even when key components are missing.
-Unlike patch-based models that rely on fixed, small regions, it extracts prototypes from semantically meaningful, wider regions.
-As a result, such methods experience sharp performance drops under occlusion due to their dependence on localized patterns, whereas PCMNet remains more robust by leveraging multiple related part concepts, preventing performance from relying too heavily on a single visual cue.
-4.5. Qualitative Results
-4.5.1. Concept Activation and Occlusion Robustness
-To further illustrate the interpretability and robustness of PCMNet, we compare the activated concepts of our method with PIPNet (Nauta et al., 2023) under clean and occluded conditions, as visualized in Fig.
-5.
+3.a shows that PCMNet degrades more gradually under occlusion compared to ProtoPNet and PIPNet.
+This robustness comes from PCMNet’s semantically rich and diverse part concepts, which capture complementary aspects of the object.
+Also, activated concepts from PCMNet preserve the faithfulness since they are extracted from not occluded regions as shown in Fig.
+3.b and buttom of Fig.
+4.
+By describing class-discriminative features from multiple perspectives, PCMNet enables more reliable predictions.
+In contrast, patch-based models depend heavily on fixed, localized cues and thus exhibit sharper performance drops.
+Qualitative Results
+Figure 4: visualizations of activated concept on test samples with matching training samples for Stanford Cars and CUB.
+To further illustrate the interpretability and robustness of PCMNet, we compare the activated concepts of our method with PIPNet under clean and occluded conditions, as visualized in Fig.
+4.
 For each method, we select two activated concepts on test images and retrieve training set examples with the highest activation values for those concepts.
@@
 Broader Concept Coverage: PCMNet extracts more diverse and semantically rich concepts compared to PIPNet.
-Our method leverages attention-based masks instead of fixed patches to capture part-level semantics over a broader spatial context, leading to more varied and interpretable regions across different object parts.
-More Informative Examples: The training examples retrieved by PCMNet for each concept exhibit more clearly activated and semantically aligned regions.
-This suggests that the discovered concepts better reflect the underlying factors influencing the model’s decision, thereby enhancing explanation quality.
-Robustness to Occlusion: Under occlusion, PIPNet often fails to highlight meaningful or consistent regions due to its reliance on limited regions.
+It leverages attention-based masks instead of fixed patches to capture part-level semantics over a broader spatial context, leading to more varied and interpretable regions across different object parts and improving explanation quality.
+Robustness to Occlusion: As shown at the bottom of Fig.
+4, under occlusion, PIPNet often fails to highlight meaningful regions due to its reliance on limited regions.
 In contrast, PCMNet maintains interpretability by activating alternative, unoccluded parts that belong to the same or related concepts.
 This demonstrates the model’s ability to generalize concept reasoning even when certain visual cues are missing.
-4.5.2. Visualization of Concept Prototypes
-To better understand the behavior of the learned prototypes in PCMNet, we visualize the extracted part-level features using t-SNE (van der Maaten and Hinton, 2008), as shown in Figure 6.
-Each point corresponds to a feature from a localized part, colors indicating the prototype (i.e., concept) and marker shapes denoting the object class.
+Figure 5: t-SNE visualization of concept-level part features extracted by PCMNet. Colors represent different prototypes, while shapes indicate object classes.
+To better understand the behavior of the learned prototypes in PCMNet, we visualize the extracted part-level features using t-SNE (tsne), as shown in Fig.
+5.
+Each point corresponds to a feature from a localized part, colors indicating the prototype (i.e.
+concept) and marker shapes denoting the object class.
 This visualization represents several insights:
-Disentanglement of Prototypes: Features assigned to different prototypes form well-separated clusters, suggesting that PCMNet effectively disentangles the latent interpretable part-based concept space.
-This separation implies that the Marginal Clustering Center loss and Concept Mining Module (CMM) encourages diverse and semantically distinct part concepts, which supports both interpretability of model decisions and robustness on occlusion.
+Disentanglement of Prototypes: Features assigned to different prototypes form well-separated clusters, indicating that PCMNet effectively disentangles part-based concepts.
+This shows that the Marginal Clustering Center loss and CMM promote diverse, semantically distinct parts, enhancing interpretability and occlusion robustness.
 Multi-Class Concept Coverage: Many prototype clusters contain the same semantic parts from multiple object classes.
@@
 For instance, different views or structural variations of a car class may be assigned to distinct concepts (Backlight vs headlight), further enhancing the model’s fine-grained interpretability based on human perception and explanation.
-5. Conclusion
+Conclusion
 We presented PCMNet, a novel framework designed to bridge the gap between accuracy and interpretability in deep learning models.
 By mining part-prototypes from dynamic image regions, PCMNet improves upon existing patch-based prototype models by ensuring semantic coherence across instances.
-Our results show that PCMNet enhances faithfulness, concept stability, and explainability, outperforming existing interpretable models such as ProtoPNet and PIPNet.
+Our results show that PCMNet enhances faithfulness, concept stability, and explainability, outperforming existing interpretable models.
 Moreover, our occlusion-based robustness analysis confirmed that PCMNet remains stable under missing or distorted input regions, demonstrating its reliability in real-world scenarios.
 Unlike conventional post-hoc explainability methods, PCMNet allows for more human-aligned, concept-based reasoning, making deep learning decisions more transparent and interpretable.
-Appendix A Additional Experiment and Results
-A.1. Implementation Details
+Acknowledgment:
+This research was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Digital Research Alliance of Canada.
+Appendix A Supplementary: Additional Experiment and Results
+Appendix A: Additional Implementation Details
+In our model, $F^{i}=\mathcal{F}(x^{i})$ denotes the spatial feature map extracted by the backbone network $\mathcal{F}$, with shape $\mathbb{R}^{D\times w\times h}$, where $D$ is the backbone feature dimension, and $w\times h$ is the spatial resolution.
+The part-prototype module $\mathcal{P}$ produces a soft assignment mask $M^{i}=\mathcal{P}(F^{i})\in\mathbb{R}^{(K+1)\times w\times h}$, where each pixel location is softly assigned to one of the $K$ part and one background channels.
+Specifically, for a given 2D location $a=(a^{x},a^{y})$, the assignment scores across all parts are:
+$M^{i}(a)=M^{i}[:,a^{x},a^{y}]\in\mathbb{R}^{K+1}.$ (9)
+while the full mask for part $p$ is given by,
+$M^{i}_{p}=M^{i}[p,:,:]\in\mathbb{R}^{w\times h}.$ (10)
 We used Resnet50 as the backbone, which extracts 2048 features.
 We project these features into 512 features before global pooling, i.e., $d_{f}=512$.
-Based on our experiment and set of hyper-parameters, for the CUBs dataset, the $d_{c}=2954$ and Stanford Cars $d_{c}=3092$ are chosen.
+Based on our experiment and set of hyper-parameters, for the CUBs dataset, the $d_{c}=2954$ and Stanford Cars $d_{c}=3092$ are computed in stage 2.
 As the DBSCAN is used for clustering inside of each class, for each part and class, the number of clusters is not fixed.
-The $\mathcal{L}_{\text{bl}}$(van der Klis et al., 2023) loss also is
-(8) $\mathcal{L}_{\text{bl}}=\mathcal{L}_{\text{cond}}+\mathcal{L}_{\text{equi}}+\mathcal{L}_{\text{pres}},$
+Learning Curvature
+Fig.
+6 illustrates the training loss and validation accuracy of PCMNet over 40 epochs.
+In Step 1, only the backbone and part-discovery modules are trained using the part discrimination loss ($\mathcal{L}\text{part}$).
+Once this loss converges, Step 2 and 3 begin, during which concept centroids are extracted and the clustering loss ($\mathcal{L}\text{cl}$) is introduced, respectively.
+This causes a temporary spike in the loss, followed by a gradual decrease as the model adapts to reasoning over learned concepts.
+To keep the prototypes dynamic and data-aligned, centroid extraction and Step 2 are repeated every 5 epochs.
+Figure 6: Training loss over 40 epochs showing three-stage learning dynamics.
+Additional details of Baseline loss
+The $\mathcal{L}_{\text{bl}}$(van2023pdisconet) loss also is
+$\mathcal{L}_{\text{bl}}=\mathcal{L}_{\text{cond}}+\mathcal{L}_{\text{equi}}+\mathcal{L}_{\text{pres}},$ (11)
 where the objectives for each loss are:
@@
   • $\mathcal{L}_{pres}$: All parts should be present at least in some of the images in the batch to avoid finding non-relevant parts.
-A.2. Additional Ablation Studies
-A.2.1. Effect of Hyper-parameters
+Appendix B: Additional Ablation Studies
+Number of Parts ($K$)
+We assess the impact of the number of parts on classification accuracy, faithfulness, and size of concept prototypes in Table 4.
+As the model extracts more prototypical concepts, its decisions become more aligned with the underlying features, enhancing interpretability.
+Interestingly, when the number of prototypes is smaller, the faithfulness scores for the top concepts tend to be higher.
+This is because the model relies on fewer, more compact concepts ($d_{c}$), concentrating its decisions on a smaller number of more influential ones.
+K Cars Birds
+C Acc F(1)-F(5) $d_{c}$ C Acc F(1)-F(5) $d_{c}$
+0 84.1 1.54 - 6.61 - 83.2 1.21-5.56 -
+4 83.5 79.96 - 91.01 1703 83.5 71.42-85.64 1786
+5 84.1 70.11 – 91.33 1979 84.9 66.7 - 85.9 2018
+6 85.2 66.02 – 90.65 2345 85.7 59.9 - 86.7 2597
+7 86.7 62.47 – 91.25 2767 86.0 55.7 - 87.3 2954
+8 87.5 59.33 – 91.73 3092 85.1 53.0 - 86.4 3242
+9 86.4 55.10 – 91.52 3277 84.4 52.5-86.1 3470
+Table 4: Impact on performance of the number of prototypes.
+Effect of Hyper-parameters.
 We investigate the sensitivity of our model to the hyper-parameters $\alpha$ and $\beta$, which control the strength of the concept consistency and interpretability regularization terms, respectively.
 As shown in Fig.
-8, our model remains robust across a wide range of values.
+7, our model remains robust across a wide range of values.
 Specifically, the classification accuracy (C Acc) and faithfulness score F(n) exhibit stable trends, with peak performance achieved around $\alpha=1.5$ and $\beta=2.0$.
@@
 The consistency of the scores across different settings highlights the reliability of the proposed regularization components.
-Figure 8.
-Impact of hyper-parameters $\alpha$ and $\beta$ on classification accuracy (C Acc) and faithfulness score F(3).
+Figure 7: Impact of hyper-parameters $\alpha$ and $\beta$ on classification accuracy (C Acc) and faithfulness score F(3).
 Our method maintains strong performance across a wide range of values, showing robustness and stability in both predictive accuracy and interpretability.
 Also, in Fig.
-9, we evaluate the impact of hyperparameters in Eq.4 on classification accuracy.
+8, we evaluate the impact of hyperparameters in Eq.4 on classification accuracy.
 At first, we set $m_{2}=1$ and measure the effect of $m_{1}$ by gradually increasing the values.
@@
 (b) $m2$
-Figure 9. The model’s classification accuracy in different values of $m1$ and $m2$.
-Figure 10.
-Visualization of activated concepts in test data and showing some examples in the training set for StanfordCars datasets.
-Each concept is shown with an index and percentage of explanation for the model decision.
-Figure 11.
-Visualization of activated concepts in test data and showing some examples in the training set for CUB-200 Birds datasets.
-Each concept is shown with an index and percentage of explanation for the model decision.
-Figure 12.
-Visualization of non-prototypical activated concepts in test data and showing some examples in the training set for (left) Stanford Cars and (right) CUB-200 Birds datasets.
-Similar to ProtoPNet and PIPNet, we select the patch for the region of the activated concept.
-A.2.2. Effect of Non-Prototypical Concepts
+Figure 8: The model’s classification accuracy in different values of $m1$ and $m2$.
+Effect of Non-Prototypical Concepts
 To evaluate the impact of different concept vector types on model performance and interpretability, we conduct an ablation study using prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors.
@@
 The full PCMNet, which combines both concept types, achieves the best results across both datasets, suggesting that the interplay between prototypical and non-prototypical representations enhances both prediction accuracy and interpretability.
+In other words, the vector $\mathbf{g}^{i}$ is used to capture features that are not assigned to any prototypical part.
+Importantly, these features do not necessarily correspond to background pixels.
+Instead, they include informative regions of the object that are not consistently detected as one of the predefined parts by the part-detection module.
+As illustrated in Fig.
+12, these non-prototypical concepts often correspond to distinctive yet less frequent features that are still valuable for decision-making.
+For example:
+  • In the Cars dataset, the “cargo” region is activated as a non-prototypical concept. Though it appears in a small subset of vehicles only, that provides strong discriminative power.
+  • In the CUB dataset, the “webbed feet” of certain birds (e.g., Albatross species) are highlighted.
+These regions are highly specific and explainable, yet do not form a consistent enough prototype across classes to be captured in the prototypical set.
 CAVs Cars Birds
@@
 PCMNet (Full) 87.5 59.3391.73 86.0 55.7 - 87.3
-Table 4.
-Impact of prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors on PCMNet’s classification accuracy (C Acc) and faithfulness scores (F(1)-F(5)) on Cars and Birds datasets.
+Table 5: Impact of prototypical ($\mathbf{z^{i}}$) and non-prototypical ($\mathbf{g^{i}}$) concept vectors on PCMNet’s classification accuracy (C Acc) and faithfulness scores (F(1)-F(5)) on Cars and Birds datasets.
 The full model achieves superior performance, highlighting the benefit of combining both concept types.
-A.3. Additional Qualitative Results
+Comparison of Performance, Interpretability, and Efficiency
+To better visualize the trade-offs among classification performance, interpretability, and computational efficiency, we present a summary plot comparing PCMNet to recent interpretable AI baselines (Fig.
+9).
+Each method is represented using a unique shape, and the color intensity of each point reflects its FLOPs (darker points indicate higher computational cost).
+The x-axis corresponds to classification accuracy on the Stanford Cars dataset, while the y-axis shows faithfulness under occlusion (F(3)), one of our main interpretability metrics.
+Figure 9: Trade-off comparison between accuracy, faithfulness (F(3)), and efficiency (FLOPs) across methods on the Stanford Cars dataset.
+Each method is represented by a distinct marker shape, and the color intensity encodes its computational cost (darker = less FLOPs).
+PCMNet achieves the best balance of high accuracy, top interpretability, and low computational cost, demonstrating its practical advantage over state-of-the-art interpretable models.
+PCMNet (full) achieves an excellent balance: it attains competitive accuracy (90.2%) and the highest interpretability (F(3) = 67.4) among all methods, while maintaining one of the lowest FLOPs (17.4G).
+In contrast, ProtoViT achieves slightly higher accuracy (91.8%) but suffers from lower interpretability (F(3) = 44.6) and significantly higher FLOPs (24.8G).
+ProtoPNet and PiPNet either underperform in accuracy or require substantially more computation.
+This visualization supports our claim that PCMNet is highly practical: it delivers state-of-the-art interpretability without sacrificing classification performance or efficiency.
+Appendix C: Additional Qualitative Results
 More results of the PCMNet prototypical concept are shown in Figs.
 10 and 11.
-12 illustrates examples of non-prototypical concepts discovered by our model for the Cars (left) and CUBs (right) datasets.
+Fig 12 illustrates examples of non-prototypical concepts discovered by our model for the Cars (left) and CUBs (right) datasets.
 For each dataset, the concept location on a test image is highlighted with a green bounding box, while corresponding activations in training examples are shown on the right using red bounding boxes.
@@
 These findings highlight the value of incorporating non-prototypical concepts in improving both interpretability and robustness of the model.
+Figure 10: Visualization of activated concepts in test data and showing some examples in the training set for StanfordCars datasets.
+Each concept is shown with an index and percentage of explanation for the model decision.
+Figure 11: Visualization of activated concepts in test data and showing some examples in the training set for CUB-200 Birds datasets.
+Each concept is shown with an index and percentage of explanation for the model decision.
+Figure 12: Visualization of non-prototypical activated concepts in test data and showing some examples in the training set for (left) Stanford Cars and (right) CUB-200 Birds datasets.
+Similar to ProtoPNet and PIPNet, we select the patch for the region of the activated concept.